### v1

# Diagnosing Vision Language Models’ Perception by Leveraging Human Methods for Color Vision Deficiencies

Abstract
Large-scale Vision Language Models (LVLMs) are increasingly being applied to a wide range of real-world multimodal applications, involving complex visual and linguistic reasoning. As these models become more integrated into practical use, they are expected to handle complex aspects of human interaction. Among these, color perception is a fundamental yet highly variable aspect of visual understanding. It differs across individuals due to biological factors such as Color Vision Deficiencies (CVDs), as well as differences in culture and language. Despite its importance, perceptual diversity has received limited attention. In our study, we evaluate LVLMs’ ability to account for individual level perceptual variation using the Ishihara Test, a widely used method for detecting CVDs. Our results show that LVLMs can explain CVDs in natural language, but they cannot simulate how people with CVDs perceive color in image based tasks. These findings highlight the need for multimodal systems that can account for color perceptual diversity and support broader discussions on perceptual inclusiveness and fairness in multimodal AI.
Diagnosing Vision Language Models’ Perception by Leveraging
Human Methods for Color Vision Deficiencies
Kazuki Hayashi        Shintaro Ozaki Yusuke Sakai        Hidetaka Kamigaito       Taro Watanabe Nara Institute of Science and Technology (NAIST), Japan {hayashi.kazuki.hl4, ozaki.shintaro.ou6 sakai.yusuke.sr9, kamigaito.h, taro.watanabe}@naist.ac.jp
1 Introduction
Large-scale Vision Language Models (LVLMs) Liu et al. (2024); Abdin et al. (2024); Ye et al. (2024); Bai et al. (2025) possess advanced reasoning capabilities that jointly process images and text, and are being integrated into real-world applications Zitkovich et al. (2023); Hao et al. (2024); Morita et al. (2024). These models should be designed with consideration for perceptual variations, ensuring accessibility and usability for individuals with differing perceptual characteristics. Color is a fundamental dimension of visual perception and a frequent reference point in everyday communication and decision making Stroop (1935). However, color perception varies across individuals due to biological and cultural differences. Color vision deficiencies (CVDs) are a common example, making it difficult to distinguish colors such as red and green Roberson and Hanley (2007); Birch (2012b).
Figure 1: A communication failure caused by color vision deficiency, where the model’s color-based directions may not be understood by some users.
Current LVLMs are trained mainly on image and text corpora that assume normal color vision Paik et al. (2021); Samin et al. (2025); Rahmanzadehgervi et al. (2024), potentially overlooking the needs of users with diverse perceptual profiles. As Figure 1 illustrates, an LVLM based navigation assistant that instructs a traveller to “Take the green line” may leave a red and green color blind user unable to act. Such mismatches degrade user experience and raise concerns about fairness and inclusivity in real-world deployments Kawakita et al. (2024).
Plate Type Answer
[Uncaptioned image] Normal 12
Protanopia 12
Deuteranopia 12
Tritanopia 12
[Uncaptioned image] Normal 8
Protanopia 3
Deuteranopia 3
Tritanopia 8
[Uncaptioned image] Normal 26
Protanopia 6
Deuteranopia 2
Tritanopia 26
Table 1: Ground-truth answers for each Ishihara plate under different vision conditions, based on clinical interpretations of the Ishihara color test Ishihara (1917).
Input Extracted mPLUG-Owl3 Output
What is Ishihara Test? … visual test used to detect color-vision deficiencies, especially red–green types such as protanopia and deuteranopia. It presents a pattern of red and green dots forming a hidden number; individuals with normal vision can read the number, whereas those with certain CVDs cannot.
What is Protanopia? color blindness that affects a person’s ability to distinguish between red and green colors. In people with protanopia, the cone cells in the retina that are responsible for perceiving red and green colors are not functioning properly.
Table 2: Representative excerpts from mPLUG-Owl3 showing detailed knowledge of the Ishihara test and CVDs. The full outputs are available in Appendix C.
Type Prompt
Base You are Protanopic.
What number do you see? Only provide the answer. No additional information.
Linguistic Support You are Protanopic.
Protanopia means you have difficulty distinguishing red tones, as red appears dimmer and may be confused with green or brown.
What number do you see? Only provide the answer. No additional information.
Visual Support You are Protanopic.
Example 1: Image shows an Ishihara plate. A Protanopic person sees this number: 2.
Now, look at the following image. What number do you see? Only provide the answer. No additional information.
Table 3: Examples of protanopia prompts. Other CVD types are provided in Appendix B.
To address this, we analyze how well LVLMs handle diverse human color perception. We focus on the Ishihara Test Ishihara (1917), a widely used and trusted diagnostic tool for assessing CVDs, and use it to diagnose the visual understanding of LVLMs under various CVD scenarios. Our experimental analysis shows that although LVLMs possess factual knowledge of CVDs, they fail to simulate and understand how people with color vision deficiencies perceive colors. This finding reveals a crucial limitation of current LVLMs in aligning with diverse human perceptual experiences and highlights the need for developing models which more fully capture diverse color perceptions.
Normal Protanopia Deuteranopia Tritanopia Avg.
LVLM Size Base Ling. Vis. Base Ling. Vis. Base Ling. Vis. Base Ling. Vis. Base Ling. Vis.
Llama-3.2 11B 14.3 9.5 5.9 5.9 11.8 5.9 4.8 23.8 9.2 11.3
LLaVA-NeXT 13B 52.4 52.4 5.9 5.9 5.9 5.9 52.4 61.9 29.2 31.5
mPLUG-Owl3 7B 71.4 71.4 62.0 5.9 5.9 5.9 5.9 5.9 17.7 66.7 66.7 62.0 37.5 37.5 36.9
Phi-3.5 4.2B 0.0 0.0 4.8 0.0 0.0 5.9 0.0 0.0 11.8 0.0 0.0 9.5 0.0 0.0 8.0
Qwen2.5-VL 7B 28.6 0.0 4.8 0.0 0.0 5.9 0.0 0.0 0.0 23.8 0.0 4.8 13.1 0.0 3.9
GPT-4o 90.5 90.5 90.5 23.6 17.7 23.6 5.9 23.5 5.9 71.4 61.9 42.9 47.9 48.4 40.7
Table 4: Accuracy (%) of each LVLM under simulated Vision types: Normal and three CVD types, with support settings: Base, Ling. (Linguistic), and Vis. (Visual). Vis. was used only for models supporting multi-image input.
2 Preliminaries
Ishihara Test
is a widely used standard tool for diagnosing red and green CVDs Ishihara (1917). It uses circular plates commonly called Ishihara plates, composed of colored dots that form numerals which appear differently depending on the viewer’s color vision type. For instance, the middle dot plates in Table 1 shows “8” under normal color vision but “3” under red and green CVDs.
Color Vision Deficiency Types Birch (2012a):
Protanopia : Reds appear darker, and it is difficult to distinguish between red, green, and brown. This is due to the absence of L-cones.
Deuteranopia : Greens may appear beige, causing red and green confusion like Protanopia but without dark reds. Caused by M-cone loss.
Tritanopia : Blues shift toward green, and yellows may appear pink or gray, making blue–yellow distinctions challenging. This is linked to the absence of S-cones.
Pilot Study.
Before evaluating whether LVLMs simulate human color perception, we first confirmed that they understand basic concepts of CVDs and the Ishihara Test. We used two prompts: (i) show an Ishihara plate and ask “What is this image used for?”, and (ii) ask for definitions such as “Please explain what Protanopia is as a type of color blindness.” As shown in Table 2, the models responded correctly, confirming readiness for the main tasks. Full outputs are in Appendix C.
Figure 2: Violin plots of mPLUG-Owl3 perplexity on Ishihara digit prediction across four vision types and three prompt settings. Lower perplexity (from force-decoding the gold answer) indicates higher model confidence.
3 Experiments and Analysis
3.1 Research Questions
We use the Ishihara Test as with humans to analyze how LVLMs perceive and simulate color vision across different task settings [^1] We use 25 out of 38 plates that require numeral responses to ensure consistent analysis across visual conditions.. Our goal is to systematically investigate LVLMs’ color perception across three analytical levels: output behavior, confidence, and internal representation.
RQ1: At the generation level, can LVLMs generate digit responses that resemble those typically reported by people with color-vision deficiencies on Ishihara plates? (§3.3)
RQ2: At the confidence level, how accurate are the model’s outputs across vision conditions, and how confident (via perplexity) is the model when producing them? (§3.4)
RQ3: At the internal representation level, what layerwise probability patterns emerge when LVLMs infer CVD types from Ishihara responses? (§3.5)
3.2 Settings
LVLMs.
We evaluate the following LVLMs: LLaVA-NeXT Liu et al. (2024), Llama-3.2 Grattafiori et al. (2024), mPLUG-Owl3 Ye et al. (2024), Phi-3.5 Abdin et al. (2024), Qwen2.5-VL Bai et al. (2025), and GPT-4o OpenAI et al. (2024). GPT-4o represents a proprietary model, and the others are chosen for their knowledge of CVDs and the Ishihara Test in pilot study of § 2. Appendix A provides more details.
Prompts.
We evaluate LVLMs using Ishihara plates and three prompt types, as shown in Table 3. The Base prompt states only the condition. The Linguistic Support prompt adds a brief explanation of how the condition affects color perception. The Visual Support prompt presents an example image showing how the plate appears to individuals with the condition, before the question.
3.3 RQ1: Generation Level
Table 4 shows model accuracy across different conditions and prompt types. Appendix D provides actual digit outputs from each model. In the Base condition, GPT-4o reaches 95 % on Normal, mPLUG-Owl3 achieves 71 %, and other models remain below 53 %. Performance collapses for both Protanopia and Deuteranopia, with no model exceeding 24 %. In contrast, Tritanopia scores exceed 65 % for GPT-4o and mPLUG-Owl3, as many plates appear visually identical to Normal. With Linguistic Support, brief descriptions shift mean accuracy by at most two points without changing rankings. GPT-4o rises to 23 % on Deuteranopia but consistently drops similarly on Tritanopia. Qwen2.5-VL often refuses when disability terms appear, averaging 0 %. Under Visual Support, example images simulating each deficiency are largely ineffective. GPT-4o drops to 41 %, Qwen2.5-VL falls by 9 %. mPLUG-Owl3 shows a modest gain on Deuteranopia, but drops on Normal and Tritanopia offset the benefit.
At the generation level, current LVLMs assume normal color vision and struggle to simulate how CVD users perceive color. Even with support, they fail to generate outputs suited to CVDs perception.
3.4 RQ2: Confidence Level
We measure token level perplexity for each forced-decoded output to evaluate how confidently LVLMs produce responses [^2] See Appendix E for details. Lower perplexity indicates higher model confidence.. Comparing perplexity across conditions reveals whether confidence aligns with accuracy, whether linguistic or visual cues reduce uncertainty, and how strongly models default to normal color vision. Figure 2 shows perplexity for mPLUG-Owl3 across four vision types and three prompts; results for other models appear in Appendix F. Perplexity remains low for Normal and Tritanopia but increases for Protanopia and Deuteranopia, suggesting that the model responds as if seeing the plates with normal color vision. Linguistic cues have little effect, and while visual support lowers perplexity for red and green deficiencies, it does not improve accuracy (Table 4).
At the confidence level, current LVLMs show consistently uncertainty in red and green tasks and fail to produce responses with appropriate confidence, even with linguistic or visual cues.
Type Prompt
Doctor-Style Inference When a person sees “{task_number}” in this Ishihara test plate, which types of color vision are possible?
Choose all that apply from the following options.
Options: - Normal - Protanopia - Deuteranopia - Tritanopia
Please output your answer.
Table 5: Doctor-style prompt, with {task_number} replaced by the correct answer for each Ishihara plate.
LVLM Size Normal Protanopia
Llama-3.2 11B 36.0 72.4
LLaVA-NeXT 13B 20.0 86.2
mPLUG-Owl3 7B 28.0 86.2
Phi-3.5 4.2B 100.0 3.45
Qwen2.5-VL 7B 4.0 86.2
Table 6: Color vision task accuracy (%) under Normal and Protanopia conditions for various LVLMs.
3.5 RQ3: Internal Representation Level
We analyze whether LVLMs can perform doctor-style diagnosis by inferring color vision types from Ishihara responses, using the instruction in Table 5. Doctor-style setting evaluates whether models can identify CVD types from response, as a doctor interprets test results. For each plate, we force-decode four labels, Normal, Tritanopia, Protanopia, Deuteranopia, and select the most probable. As shown in Table 6, models perform well on Protanopia and Deuteranopia (over 70 %, except Phi-3.5), but poorly on Normal and Tritanopia (under 30 %). This contrasts with Table 4, where they failed to recognize digits under red and green conditions.
Figure 3: LLaVA LogitLens averages. Left: Normal/Tritanopia; Right: Protanopia/Deuteranopia.
To further examine this behavior, we apply Logit Lens nostalgebraist (2020) to analyze token level probabilities across layers, following prior work showing that intermediate signals can influence outputs Wendler et al. (2024); Schut et al. (2025). Details of this computation appear in Appendix G. Figure 3 shows average diagnosis probabilities in LLaVA-NeXT. Results for other models are in Appendix H. LLaVA-NeXT tends to interpret images as Normal in mid-layers, shifting to Protanopia at the final layer, suggesting inconsistent processing. Nearly identical curves across ground-truths imply poor differentiation between conditions.
At the internal representation level, current LVLMs do not clearly distinguish color vision types. They rely on surface cues like readable digits, rather than using perceptual reasoning.
4 Conclusion
In this study, we evaluated whether LVLMs can handle diverse aspects of human color perception, especially CVDs, by using the widely adopted Ishihara Test as a benchmark for perceptual simulation. We systematically analyzed this capability by examining generation, confidence, and internal representation. The results show that although current LVLMs have accurate linguistic knowledge about CVDs, they do not understand how CVD users perceive color. They generate inappropriate outputs, exhibit low confidence in red and green tasks, and lack internal representations that distinguish vision types, indicating a limited consideration of perceptual aspects of color. These findings highlight the need to design LVLMs that go beyond text descriptions and account for perceptual differences, making AI more accessible and user-friendly.
5 Limitations
Dataset scope and scale.
Our evaluation relies on a small set of Ishihara plates that primarily target red and green color-vision deficiencies. The limited number and narrow focus reduce the diversity of conditions tested, so our findings may not generalise to other CVD types or broader visual tasks. Nevertheless, this short paper fulfills its objective by demonstrating the importance of analyzing model behavior in depth, specifically from the perspective of color perception, as a step toward building user-friendly and perceptually inclusive LVLMs. While scaling to broader datasets and real-world tasks remains important future work, our focus here is on establishing a methodology and highlighting key limitations in current models.
Limitations in Evaluating Tritanopia.
The Ishihara Test is primarily designed to detect red and green deficiencies (Protanopia and Deuteranopia), and is not suitable for evaluating blue–yellow deficiencies such as Tritanopia. As such, our findings may not fully capture how LVLMs handle Tritanopia cases. Nevertheless, we included Tritanopia-specific instructions to test whether models follow the prompt condition or respond based on the visible digit. Future work should evaluate all major CVD types, including Tritanopia, using more general tasks beyond Ishihara plates.
Evaluation methodology.
We prompt LVLMs to simulate color-blind perception in a controlled, diagnostic task (identifying digits in Ishihara patterns). This single evaluation format may not capture the full complexity of color reasoning. Prior studies show that changing task formulations alone can yield different outcomes. Moreover, we assume the Ishihara ground-truth digits as an oracle for correct behavior, without validating outputs with people who have CVD or ophthalmologists.
Generalisability of model behavior.
Success or failure on stylised Ishihara plates does not guarantee that a model truly perceives color differences across diverse, real-world settings. An LVLM might exploit dot-pattern heuristics or memorised cues rather than authentically experiencing color confusion. Prior work shows vision–language models can be biased toward certain colors and struggle with others Raj et al. (2024), suggesting performance in one controlled test may not translate to broad color understanding. Caution is therefore needed when extrapolating our results to natural images with varied palettes or complex contexts.
6 Ethical Considerations
Inclusive design for perceptual diversity.
Recent AI research communities, including ACL, have emphasized the importance of designing fair and inclusive systems that accommodate users with perceptual or cognitive differences. Color-vision deficiency, which affects a significant portion of the global population, is one such factor that is often overlooked in model development and evaluation. By explicitly testing LVLM behavior under CVD conditions, this study contributes to a broader understanding of accessibility in multimodal AI systems and encourages further research in this direction. Building on this motivation, we also discuss fairness-related concerns in the context of model evaluation.
Not for Diagnostic Use.
This study aims to evaluate whether models can simulate aspects of color vision deficiency as a step toward developing systems that better capture the diversity of human perception. It is not intended to replace professional diagnosis or support clinical decision-making. Even if a model produces correct responses to certain vision tests, these outputs must not be used for medical purposes or institutional assessments. To prevent such misuse, we clearly define the scope and limitations of our evaluation and explicitly state that the outputs are not a substitute for expert judgment in clinical contexts.
Licensing and copyright.
The Ishihara Test Ishihara (1917) is a long-established diagnostic tool widely used in academic research across fields such as ophthalmology, psychology, education, and machine learning (Zhao et al., 2019; Klinke et al., 2024; Lin et al., 2019). We did not reproduce, modify, or redistribute the plates; they were used solely for research and analysis.
AI-assistant usage.
Portions of the manuscript, such as prompt templates and wording adjustments, were drafted with the assistance of GPT-4 to ensure linguistic clarity. All technical content, analysis scripts, and final decisions were made by the research team.
Appendix A Detailed Model Settings
In this study, to ensure fair and consistent performance comparisons, all experiments were conducted using a single NVIDIA RTX 6000 Ada GPU. All models were run under the same computational settings, with generation performed using half quantization to reduce memory usage while maintaining model fidelity. We used publicly available models from Hugging Face and OpenAI, as summarized in Table 7.
Model Params HuggingFace Name / OpenAI Name
Phi3.5 4.2B microsoft/Phi-3-vision-128k-instruct
Qwen2.5-VL 7B Qwen/Qwen-2.5-VL-7B-Instruc
mPLUG 7B mPLUG/mPLUG-Owl3-7B-240728
Llama3.2 11B meta-llama/Llama-3.2-11B-Vision-Instruct
LLaVA-NeXT 13B llava-hf/llava-v1.6-vicuna-13b-hf
GPT-4o gpt-4o-2024-11-20
Table 7: Detailed model names.
Appendix B Prompt Details
To evaluate how LVLMs respond under different levels of guidance, we design three prompt types for each color vision condition: Base, Linguistic Support, and Visual Support. Each targets a distinct aspect of the model’s capacity to simulate color perception under CVDs.
Base Prompts
Base prompts provide only the color vision condition, for example “You are Protanopic”, and ask the model to report the number shown in the Ishihara plate. This setup tests whether the model can simulate CVDs using only the condition label. See Table 8.
Prompts with Linguistic Support
These prompts add a textual description of the impairment, such as noting that red tones appear darker. This setting assesses whether linguistic context improves predictions. See Table 10.
Prompts with Visual Support
These prompts include a brief condition description and a reference example showing what number appear in another plate. The model is then asked to identify the number in a new image. This tests whether few-shot visual examples improve simulation. See Table 9. All prompts end with: “Only provide the answer. If you see nothing, answer ‘Nothing’. No additional information.” This standardization allows consistent, quantitative comparison across all settings.
Appendix C Language Knowledge of the Ishihara Test and CVDs
Before testing visual ability, we verified that all LVLMs possess basic language knowledge of the Ishihara Test and the three major color vision deficiencies. As the main paper includes only one example (Table 2), this appendix provides full outputs for the remaining models.
Tables 1418 show representative answers from Llama-3.2, LLaVA-NeXT, mPLUG-Owl3, Phi-3.5, and Qwen2.5-VL. Each table has input prompts on the left and trimmed model responses on the right. All models—except Phi-3.5, which returned an error for the Ishihara query—accurately describe the test and define Protanopia, Deuteranopia, and Tritanopia, confirming sufficient textual knowledge. These results show that failures in the main experiments arise from weaknesses in visual grounding, not language understanding.
Appendix D Plate-level Model Outputs
The main paper reports only aggregated accuracy due to space constraints. Since the Ishihara subset includes just 25 plates, we provide the complete set of raw predictions here. Tables 11, 12, and 13 show the digits predicted by each model under the Base, Linguistic Support, and Visual Support prompts. The first row labeled Gold indicates the reference digit for each condition [^3] Plates without a digit or with “Nothing” as the correct response are marked as N/A.. Lower rows show the outputs returned by each LVLM.
Base prompt outputs
Table 11 shows model predictions and gold digits. In Protanopia and Deuteranopia, most models return digits matching the Normal condition, indicating failure to simulate red and green confusion. Failure patterns vary: Phi-3.5 produces no outputs, while Qwen2.5-VL repeatedly generates "70" or similar values, likely due to decoding issues. GPT-4o and mPLUG-Owl3 perform well on Normal and Tritanopia but diverge under red and green conditions. These results show that strong performance under normal vision does not imply accurate CVD simulation. The Base setting shows that current LVLMs often rely on shallow heuristics and fail to capture red and green perceptual variation, as reflected in Table 4.
Linguistic support outputs
Table 12 shows predictions under prompts that include brief descriptions of each condition. Despite the added context, accuracy remains similar to the Base setting. Models often respond as if assuming normal vision, suggesting that descriptive text does not influence internal visual reasoning. Qwen2.5-VL outputs under Base but fails under Linguistic Support, suggesting that disability-related terms can disrupt generation. Overall, text-only prompts neither improve grounding nor help simulate perceptual differences, showing the limitations of linguistic cues alone.
Visual support outputs
Table 13 shows predictions under Visual Support prompts, which include a reference image before the target plate. LLaVA-NeXT and Llama-3.2 are excluded due to lack of multi-image support. Evaluated models are mPLUG-Owl3, Phi-3.5, Qwen2.5-VL, and GPT-4o. No model improves over the Base setting, and many perform worse. mPLUG-Owl3 and GPT-4o often fail on plates they previously answered correctly, suggesting the added image introduces noise rather than clarity. N/A responses increase, with GPT-4o and Qwen2.5-VL more often producing empty or degenerate outputs. These results suggest that visual examples may cause confusion or overgeneralization. Overall, Visual Support fails to enhance simulation and often increases errors, showing that current LVLMs struggle to leverage visual demonstrations effectively.
Appendix E Per-token Perplexity Computation
We report here the exact procedure used to derive the confidence scores in §3.4. For each Ishihara plate, the image, prompt, and gold numeral answer are concatenated into a single input sequence. All tokens that precede the answer are masked with ignore_index, ensuring that only the answer tokens contribute to the loss. Given answer tokens $x_{1\!\,:\,\!L}$, we obtain the token-level log-probabilities:
$v_{i}=\log P(x_{i}\mid\texttt{prompt},x_{1\!\,:\,\!i-1}),$
and compute the average negative log-likelihood:
$\mathcal{L}=-\frac{1}{L}\sum_{i=1}^{L}v_{i}\quad\text{and}\quad\mathrm{PPL}=\exp(\mathcal{L}).$
Lower $\mathrm{PPL}$ thus corresponds to higher model confidence in the forced-decoded answer. All perplexity figures reported in Figure 2 and Appendix F are obtained via this definition.
Appendix F Perplexity and Model Confidence
Token-level perplexity is measured for Llama-3.2, mPLUG-Owl3, Phi-3.5, and Qwen2.5-VL. LLaVA-NeXT hides the answer during forced decoding, and GPT-4o does not return token probabilities, so both are excluded.
Base prompt
Table 19 reports mean and standard deviation of perplexity across vision types. All four models show the same trend: very low values for Normal and Tritanopia, and much higher ones for red and green deficiencies. For instance, Llama-3.2 rises from 33716.2 to 513.0, and mPLUG-Owl3 from 5.3 to 353.0. This large gap, along with low variance for Normal and high for CVDs, suggests a strong bias toward normal vision.
Linguistic Support
Table 20 shows that brief textual descriptions have little effect on the mean and often increase variance. Llama-3.2 and Phi-3.5 widen the gap between Normal and CVDs. Phi-3.5 rises from 17.4 to 171.7, while Qwen2.5-VL changes only slightly. Thus, linguistic context fails to shift the models’ perceptual prior.
Visual Support
Table 21 covers three models with multi-image input. mPLUG-Owl3 drops from 390.2 to 39.4, and Phi-3.5 from 171.7 to 11.1, but neither improves accuracy. Qwen2.5-VL becomes unstable, with Protanopia variance at 362.0. Little change in means and higher variance suggest visual support adds noise, not clarity.
Appendix G Logit Lens Probability Computation
At each transformer layer we project the hidden state $h$ back into the vocabulary space with the unembedding matrix $W_{\text{unembed}}$ and apply a softmax:
$\text{LogitLens}(h)\;=\;\text{softmax}\bigl{(}h\cdot W_{\text{unembed}}\bigr{)}.$
The resulting distribution gives the probability of every token at that layer. For a specific diagnosis-label token $t$ (e.g., Protanopia or Deuteranopia) we use
$P(t)\;=\;\frac{\exp\!\bigl{(}(h\cdot W_{\text{unembed}})_{t}\bigr{)}}{\sum_{j}\exp\!\bigl{(}(h\cdot W_{\text{unembed}})_{j}\bigr{)}}.$
Here, $(h\cdot W_{\text{unembed}})_{t}$ denotes the logit for token $t$, and the denominator normalizes over all vocabulary tokens. This $P(t)$ is the layer-wise probability assigned to the diagnosis label.
Appendix H Detailed Diagnosis of CVDs with LVLMs
Figures 710 extend the main analysis by visualising token-level label probabilities across transformer layers for four additional LVLMs. Each panel groups plates by the correct vision condition and plots plate-wise trajectories (thin lines) together with the average curve (bold line). Left plots correspond to Normal or Tritanopia, right plots to Protanopia or Deuteranopia.
Llama-3.2 (Figure 7), For Normal/Tritanopia the Normal label stays near zero until three layers before the head, then rises above 50%. red and green show an analogous late surge in the Protanopia curve, mirroring the perplexity jump in Table 19.
mPLUG-Owl3 (Figure 8), Early layers are mixed but converge rapidly. Tritanopia exceed 80% by layer 26. When the digit is unreadable, the model shifts probability toward Tritanopia, possibly explaining its low variance under Base.
Phi-3.5 (Figure 9), Curves remain diffuse and never exceed 25% even near the output. This matches the wide perplexity spread and frequent no-answer cases in Table 11.
Qwen2.5-VL (Figure 10), Normal show local peaks rather than a smooth ascent. red and green plates present competing red and green waves that collapse to a single token at the final layer. However, the overall probabilities remain extremely low, on the order of $10^{-6}$ even beyond layer 6, which indicates very weak model confidence.
Condition Prompt
Protanopia You are Protanopic.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Deuteranopia You are Deuteranopic.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Tritanopia You are Tritanopic.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Normal Vision What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Table 8: Base prompts for each color-vision condition.
Condition Prompt
Protanopia You are Protanopic.
You have difficulty distinguishing certain colors.
Example 1:
Image shows an Ishihara plate.
A Protanopic person sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Deuteranopia You are Deuteranopic.
You have difficulty distinguishing certain colors.
Example 1:
Image shows an Ishihara plate.
A Deuteranopic person sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Tritanopia You are Tritanopic.
You have difficulty distinguishing certain colors.
Example 1:
Image shows an Ishihara plate.
A Tritanopic person sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Normal Vision You have normal color vision.
Example 1:
Image shows an Ishihara plate.
A person with normal color vision sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Table 9: Prompts with visual support examples for each color-vision condition.
Setting Prompt
Protanopia You are Protanopic.
Protanopia means you have difficulty distinguishing red tones, as red appears dimmer and may be confused with green or brown.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Deuteranopia You are Deuteranopic.
Deuteranopia means you have difficulty distinguishing green tones, as green appears dimmer and may be confused with red or brown.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Tritanopia You are Tritanopic.
Tritanopia means you have difficulty distinguishing blue tones, as blue appears dimmer and may be confused with green or gray.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Normal Vision What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Table 10: Prompts with linguistic support for each color-vision condition.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
Llama 19 8 8 42 7 1 2 4 15 2 42 42 1 7 7 4 53 15 15 3 4 52 52 52 8
Llava 12 8 6 29 37 3 6 16 7 2 6 9 5 6 7 16 3 1 1 1 1 28 42 35 96
mPlug 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 100 100 10 100 23 42 35 96
Phi
Qwen 12 8 6 29 57 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 48 74 95
GPT 12 8 6 29 57 5 3 15 74 2 6 74 45 5 7 16 73 74 74 74 5 93 42 35 96
Protanopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 6 2 5 6
Llama 19 2 5 17 3 3 2 6 3 2 2 2 45 7 4 6 46 73 4 53 52 4 4 4 2
Llava 12 8 6 29 57 3 6 16 7 2 6 9 1 6 7 16 3 1 1 1 1 20 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 1 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 74 7 74 74 74 74 38 43 74 95
GPT 12 3 3 29 55 3 5 15 4 9 3 37 45 N/A 7 16 13 74 2 74 27 25 N/A 39 96
Deuteranopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 2 4 3 9
Llama 19 2 5 52 1 53 52 49 45 53 52 52 36 55 7 56 4 4 2 49 53 42 52 37 8
Llava 12 8 6 29 57 3 3 16 14 2 6 9 10 6 7 16 3 10 1 1 12 20 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 1 1 10 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 53 7 74 74 74 74 88 43 30 95
GPT 12 8 6 29 57 3 8 N/A 74 2 6 97 45 3 7 16 73 74 N/A N/A 5 93 42 35 96
Tritanopia Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
Llama 19 2 5 17 1 3 2 4 1 3 4 42 46 4 7 3 8 4 4 4 1 3 4 3 8
Llava 12 8 6 29 37 3 3 16 7 2 6 9 1 6 7 16 3 1 1 1 12 28 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 10 10 10 10 23 42 35 96
Phi
Qwen 21 8 6 29 37 74 74 74 74 74 74 74 74 5 3 53 3 74 0 74 74 38 43 30 95
GPT 12 8 6 29 57 N/A 3 15 74 2 3 37 36 5 7 16 33 74 74 6 3 N/A 42 35 96
Table 11: Predicted outputs of six LVLMs (Llama, Llava, mPlug, Phi, Qwen, and GPT) on Ishihara color-vision test plates under four simulated conditions (Normal, Protanopia, Deuteranopia, Tritanopia) using the Base prompt only. The “Gold” row shows the ground-truth digits.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
LLama 16 8 8 42 52 4 4 52 15 3 15 42 3 4 7 4 15 15 4 15 6 3 4 52 52
LLava 12 8 6 29 37 3 6 16 7 2 6 9 5 6 7 16 3 1 1 1 1 28 42 35 96
mPlug 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 100 100 10 100 23 42 35 96
Phi
Qwen
GPT 12 8 6 29 57 5 3 15 74 2 6 74 45 5 7 16 23 26 74 29 5 26 42 35 96
Protanopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 6 2 5 6
LLama 15 8 7 42 5 6 2 3 9 8 4 2 3 4 4 3 6 2 6 4 4 6 52 2 42
LLava 12 8 6 29 37 3 6 16 7 2 6 9 1 6 7 16 3 1 1 1 1 26 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 9 1 23 42 35 96
Phi
Qwen
GPT 32 3 5 26 5 8 8 13 21 5 3 37 46 2 7 16 13 3 5 5 5 23 12 30 39
Deuteranopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 2 4 3 9
LLama 8 8 6 59 3 4 6 4 66 6 8 22 49 7 7 4 4 8 7 11 41 2 42 33 44
LLava 12 8 6 29 57 3 6 16 14 2 6 9 1 6 7 16 3 1 1 12 12 26 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 9 1 23 42 35 96
Phi
Qwen
GPT 12 3 5 29 37 2 8 13 41 2 5 97 45 3 3 16 78 6 5 3 74 93 42 35 96
Tritanopia Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
LLama 12 2 8 17 8 4 3 6 2 2 2 8 7 5 7 7 6 2 4 2 4 6 4 4 8
LLava 12 8 6 29 57 3 3 16 7 2 6 9 1 6 7 16 3 1 1 12 12 26 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 9 1 23 42 35 96
Phi
Qwen
GPT 12 8 6 29 57 2 8 15 41 2 3 73 35 5 7 16 73 3 5 5 3 93 42 35 96
Table 12: Predicted outputs of six LVLMs (Llama, Llava, mPlug, Phi, Qwen, and GPT) on Ishihara color-vision test plates under four simulated conditions, using Linguistic Support prompts that include a brief explanation of how each deficiency alters color perception. The “Gold” row shows the ground-truth digits.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
mPlug 12 8 6 29 37 5 3 15 2 6 97 90 5 5 7 16 13 N/A N/A N/A 26 23 42 35 96
Phi 12 6 29 37 5 3 15 74 2 6 97 45 5 7 16 73 1 3 3 3 25 42 3 96 90
Qwen 8 6 38 37 74 74 15 74 74 74 3 37 74 3 16 73 4 74 42 74 74 37 24 74 74
GPT 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 33 74 74 N/A 74 93 42 35 96
Protanopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 6 2 5 6
mPlug 12 8 6 29 37 5 3 15 2 2 6 9 5 5 7 13 13 1 45 73 6 2 43 3 96
Phi 12 6 6 35 3 1 17 1 3 3 4 17 3 7 7 3 1 2 45 73 3 3 3 3 9
Qwen 7 3 8 74 7 7 3 74 N/A 74 74 74 3 3 8 N/A 7 7 74 42 4 7 3 3 7
GPT 21 3 5 26 5 5 5 15 N/A 2 3 37 45 5 N/A 16 3 5 74 6 3 25 42 33 96
Deuteranopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 2 4 3 9
mPlug 12 8 6 29 37 5 3 15 2 2 6 9 5 5 7 13 13 1 45 73 2 2 4 35 96
Phi 12 5 35 17 7 7 45 73 3 3 9 9
Qwen 7 7 3 74 7 7 23 74 74 74 74 74 74 74 8 74 7 7 74 42 7 7 7 7 7
GPT 14 0 5 29 37 5 9 15 74 2 3 97 45 5 N/A 16 33 2 17 74 5 23 42 35 96
Tritanopia Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
mPlug 12 8 6 29 37 5 3 15 2 6 6 90 5 5 7 13 3 N/A N/A N/A 26 42 42 35 96
Phi 12 6 23 5 3 15 74 6 9 45 7 7 73 26 42 3 96 9
Qwen 3 3 3 3 4 7 15 3 4 3 3 3 3 3 1 42 N/A N/A N/A N/A 4 3 42 3 3
GPT 21 3 6 29 N/A 5 8 15 71 2 5 57 46 3 3 16 N/A N/A N/A N/A 57 26 42 35 86
Table 13: Predicted outputs of four LVLMs (mPlug, Phi, Qwen, and GPT) on Ishihara color-vision test plates under four simulated conditions, using Visual Support prompts that present an illustrative example image of each plate as seen under the deficiency. The “Gold” row shows the ground-truth digits. Evaluation was performed only on models capable of ingesting two or more images in a single turn.
Input Model Output (LLama as an example)
What is Ishihara Test? This image is used to test color vision. It is called an Ishihara plate, named after its creator, Dr. Shinobu Ishihara, a Japanese professor of ophthalmology. The plate is a widely used tool for diagnosing color blindness, particularly red-green color blindness, which is the most common form of color vision deficiency. The plate consists of a series of dots in different colors and sizes, arranged in a specific pattern. The viewer is asked to identify a number or shape hidden within the dots. For people with normal color vision, the correct answer is obvious, but for those with color vision deficiency, the correct answer may not be apparent. In this image, the number 8̈ïs hidden among the dots. The dots are arranged in a way that creates a red 8̈s̈hape, but the dots are also scattered randomly throughout the plate, making it difficult to see the number without looking closely. For someone with normal color vision, the red 8̈s̈tands out clearly, but for someone with red-green color blindness, the number may not be visible at all. The Ishihara plate is used to diagnose color blindness because it exploits the fact that people with certain types of color vision deficiency have difficulty distinguishing between certain colors. In this case, the plate uses the fact that people with red-green color blindness have trouble distinguishing between red and green colors. By using a combination of red and green dots, the plate can help identify individuals who have difficulty seeing these colors. Overall, the Ishihara plate is a simple yet effective tool for diagnosing color blindness, and it has been widely used in medical settings for many years.
What is Protanopia? Protanopia, also known as red-blindness, is a type of color vision deficiency that affects approximately 1% of males and 0.05% of females. It is caused by a mutation in the OPN1LW gene, which codes for the long-wavelength cone cells in the retina. These cone cells are responsible for detecting long-wavelength light, which is associated with the color red. In individuals with protanopia, the long-wavelength cone cells are either absent or non-functional, leading to a reduced ability to detect red light. As a result, people with protanopia may have difficulty distinguishing between red
What is Deuteranopia? Deuteranopia is a type of color vision deficiency that affects approximately 5% of males and 0.4% of females in the United States. Deuteranopia, also known as red-green color blindness, is a type of color vision deficiency that affects a person’s ability to see certain colors, particularly red and green. People with deuteranopia have difficulty distinguishing between these two colors and may see them as similar shades of yellow or brown. Deuteranopia is caused by a deficiency in the long-wavelength cones in the retina, which are responsible for detecting red light. As a result, people with deuter
What is Tritanopia? Tritanopia is a rare form of color vision deficiency that affects a person’s ability to see the color blue. People with tritanopia have difficulty distinguishing between different shades of blue and green, and may see these colors as more yellowish or brownish. Tritanopia is a congenital condition, meaning it is present at birth, and is usually inherited in an X-linked recessive pattern. This means that the gene responsible for tritanopia is located on the X chromosome, and males are more likely to be affected because they have only one X chromosome. Females, on the other hand, are less likely to be affected
Table 14: An example of Llama’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( LLava as an example)
What is Ishihara Test? This image appears to be a color vision test, also known as an Ishihara test. The test is used to check for the presence of a condition called central serous chorioretinopathy (CSCR), which can cause central serous chorioretinopathy. The test uses a pattern of colored dots, and the presence of certain colors can indicate the presence of the condition. The pattern is typically circular and contains a number of colored dots. If a person can see a specific color pattern, it may suggest the presence of CSCR. This test is commonly used by ophthalmologists to diagnose this condition.
What is Protanopia? Protanopia is a type of color vision deficiency, also known as color blindness. It is a genetic condition that affects the way an individual perceives colors. People with protanopia have difficulty distinguishing between red and green hues. In protanopia, the red photopigment in the retina is either missing or not functioning properly. As a result, individuals with protanopia may see red as a shade of brown or green, and they may have difficulty distinguishing between different shades of red and green. Protanopia
What is Deuteranopia? Deuteranomaly, also known as deuteranomaly or deuteranomaly, is a type of color blindness that affects the perception of colors. It is caused by a deficiency in one of the three types of cone cells in the retina of the eye, which are responsible for detecting different wavelengths of light. People with deuteranomaly have difficulty distinguishing between certain colors, such as red and green, or blue and yellow. The severity of deuteranomaly can vary from mild to severe, and it can affect
What is Tritanopia? Tritanopia is a type of color vision deficiency, also known as color blindness. It affects the perception of colors, making it difficult for individuals with this condition to distinguish between certain colors. In Tritanopia, individuals have difficulty distinguishing between the colors blue and green. This is because the red and green photopigments in the cone cells of the retina, which are responsible for detecting these colors, are less sensitive or absent in individuals with Tritanopia. As a result, individuals with Tritanopia may see blue and
Table 15: An example of LLava’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( mPlug as an example)
What is Ishihara Test? The image is a visual test known as the Ishihara test, which is used to determine whether a person has protanopia, a type of color blindness. The test works by presenting a pattern of dots in various colors, and the individual is asked to identify the number of dots in the pattern. People with protanopia may have difficulty distinguishing between red and green colors, which is why the number 6 is used in this test. The test is designed to be non-invasive and can be used to identify color vision deficiencies without the need for medical equipment.
What is Protanopia? of color blindness that affects a person’s ability to distinguish between red and green colors. It is the most common form of color blindness, affecting approximately 8% of men and 0.5% of women. In people with protanopia, the cone cells in the retina that are responsible for perceiving red and green colors are not functioning properly. This can result in difficulty distinguishing between red and green, as well as other shades of color. In some cases, people with protanopia may see red and green colors as different shades of yellow or blue.
What is Deuteranopia? of color vision deficiency that affects a person’s ability to distinguish between red and green colors. It is the most common form of color blindness, affecting approximately 3% of the male population and 0.5% of the female population worldwide. People with deuteranopia have difficulty distinguishing between the red and green wavelengths of light. This is because the cone cells in their eyes, which are responsible for color vision, are not functioning properly. Specifically, the cone cells that are responsible for detecting red and green wavelengths are not working as well as they should be
What is Tritanopia? to distinguish between blue and green colors. It is caused by a deficiency in one of the three types of cone cells in the retina of the eye, which are responsible for detecting different wavelengths of light. In people with tritanopia, the cone cells that are responsible for detecting blue light are not functioning properly, which can make it difficult for them to distinguish between blue and green colors. This type of color blindness is relatively rare, affecting only a small percentage of the population.
Table 16: An example of mPlug’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( Phi as an example)
What is Ishihara Test? Error generating response.
What is Protanopia? Protanopia, also known as total red color blindness, is a type of color blindness where individuals have difficulty distinguishing between red and green hues. This condition is caused by a genetic mutation that affects the cone cells in the retina, which are responsible for detecting red light. As a result, people with protanopia see the world in shades of gray, green, and blue, and they cannot perceive the full spectrum of colors that most people can see. Instruction 1 (Same difficulty/format/length):
What is Deuteranopia? Deuteranopia is a form of color blindness that affects an individual’s ability to distinguish between red and green hues. It is a hereditary condition caused by the absence or malfunction of the green cone cells in the retina, which are responsible for detecting green light. People with deuteranopia have difficulty with tasks that require color discrimination, such as reading color-coded maps or distinguishing ripe fruits from unripe ones. The severity of deuteranopia can vary from mild to severe, with some individuals being able to perceive
What is Tritanopia? Tritanopia is a form of color blindness where individuals have difficulty distinguishing between blue and yellow hues. This condition is a type of anomalous trichromacy, which means that one of the three cone cells in the eye responsible for color vision is not functioning properly. In the case of tritanopia, the blue cone cells are affected, leading to a reduced ability to perceive blues and yellows. People with tritanopia often see these colors as shades of yellow or green. This condition is less common than red-green color blindness and affects approximately
Table 17: An example of Phi’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( Qwen as an example)
What is Ishihara Test? This image is an Ishihara color vision test plate, commonly used to assess color blindness, particularly red-green color vision deficiencies. The test consists of a series of plates with colored dots that form numbers or letters when viewed by someone with normal color vision
What is Protanopia? 1. Protanopia is a type of color blindness that affects a person’s ability to distinguish between red and green colors. 2. It is caused by a genetic mutation that affects the cone cells in the retina of the eye, which are responsible for detecting colors. 3. People with protanopia may have difficulty distinguishing between shades of red and green, and may see these colors as more similar to each other than they actually are. 4. The severity of protanopia can vary from person to person, with some individuals having a mild form of the condition and others having a more severe form. 5. There is
What is Deuteranopia? 1. Deuteranopia is a type of color blindness that affects the ability to distinguish between red and green colors. 2. People with deuteranopia have difficulty distinguishing between shades of red and green, and may see these colors as more similar than they actually are. 3. Deuteranopia is caused by a genetic mutation that affects the cone cells in the retina of the eye, which are responsible for detecting color. 4. It is estimated that deuteranopia affects approximately 1 in 12 men and 1 in 200 women. 5. People with de
What is Tritanopia? 1. Tritanopia is a type of color blindness that affects the perception of blue and green colors. 2. People with Tritanopia have difficulty distinguishing between blue and green hues, and may see these colors as more similar to each other than they actually are. 3. Tritanopia is caused by a genetic mutation that affects the cone cells in the retina, which are responsible for detecting blue and green light. 4. There are two main types of Tritanopia: Tritanomaly and Tritanopia. Tritanomaly is a milder form of the condition, while Tritanopia is more severe
Table 18: An example of Qwen’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Figure 4: Violin plots of LLama perplexity distributions for Ishihara digit predictions under four simulated vision conditions (Normal–blue, Protanopia–red, Deuteranopia–green, Tritanopia–orange) and three prompt types (Base, Linguistic, Visual). For each condition, perplexity is measured by force decoding the Gold Answer corresponding to the simulated vision type. The plots show how prompt context and vision condition affect model confidence.
Figure 5: Violin plots of Phi perplexity distributions for Ishihara digit predictions under four simulated vision conditions (Normal–blue, Protanopia–red, Deuteranopia–green, Tritanopia–orange) and three prompt types (Base, Linguistic, Visual). For each condition, perplexity is measured by force decoding the Gold Answer corresponding to the simulated vision type. The plots show how prompt context and vision condition affect model confidence.
Figure 6: Violin plots of Qwen perplexity distributions for Ishihara digit predictions under four simulated vision conditions (Normal–blue, Protanopia–red, Deuteranopia–green, Tritanopia–orange) and three prompt types (Base, Linguistic, Visual). For each condition, perplexity is measured by force decoding the Gold Answer corresponding to the simulated vision type. The plots show how prompt context and vision condition affect model confidence.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size Mean SD Mean SD Mean SD Mean SD
LLaMA 11B 33716.2 176261.9 513.0 1453.1 524.7 1451.5 33716.2 176261.9
mPLUG-Owl3 7B 5.3 7.7 353.0 646.4 331.9 652.0 5.3 7.7
Phi-3.5 4.2B 17.2 25.3 119.5 406.4 88.6 404.2 17.2 25.3
Qwen2.5-VL 7B 31.4 62.2 64.5 91.2 44.9 74.4 31.4 62.2
Table 19: Mean and standard deviation (SD) of perplexity for each vision condition under the Base prompt setting.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size Mean SD Mean SD Mean SD Mean SD
LLaMA 11B 14140.3 66220.6 596.8 1525.2 605.8 1523.3 14140.3 66220.6
mPLUG-Owl3 7B 5.3 7.4 309.2 621.3 288.0 626.2 5.3 7.4
Phi-3.5 4.2B 17.4 22.9 171.7 531.1 116.4 522.7 17.4 22.9
Qwen2.5-VL 7B 32.0 59.6 58.0 85.4 41.3 69.8 32.0 59.6
Table 20: Mean and standard deviation (SD) of perplexity for each vision condition under the Linguistic prompt setting.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size Mean SD Mean SD Mean SD Mean SD
LLaMA 11B
mPLUG-Owl3 7B 11.0 27.3 39.4 55.7 30.8 47.3 11.0 27.3
Phi-3.5 4.2B 8.9 11.3 11.1 13.2 9.9 13.3 8.9 11.3
Qwen2.5-VL 7B 100.7 132.9 248.3 362.0 149.8 203.0 100.7 132.9
Table 21: Mean and standard deviation (SD) of perplexity for each vision condition under the Visual prompt setting.
Figure 7: In the Llama model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
Figure 8: In the mPlug model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
Figure 9: In the Phi model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
Figure 10: In the Qwen model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.