### v2

# Diagnosing Vision Language Models’ Perception by Leveraging Human Methods for Color Vision Deficiencies

Abstract
Large-scale Vision-Language Models (LVLMs) are being deployed in real-world settings that require visual inference. As capabilities improve, applications in navigation, education, and accessibility are becoming practical. These settings require accommodation of perceptual variation rather than assuming a uniform visual experience. Color perception illustrates this requirement: it is central to visual understanding yet varies across individuals due to Color Vision Deficiencies, an aspect largely ignored in multimodal AI. In this work, we examine whether LVLMs can account for variation in color perception using the Ishihara Test. We evaluate model behavior through generation, confidence, and internal representation, using Ishihara plates as controlled stimuli that expose perceptual differences. Although models possess factual knowledge about color vision deficiencies and can describe the test, they fail to reproduce the perceptual outcomes experienced by affected individuals and instead default to normative color perception. These results indicate that current systems lack mechanisms for representing alternative perceptual experiences, raising concerns for accessibility and inclusive deployment in multimodal settings.
[…]
Diagnosing Vision Language Models’ Perception by Leveraging
Human Methods for Color Vision Deficiencies
Kazuki Hayashi    Shintaro Ozaki Yusuke Sakai    Hidetaka Kamigaito    Taro Watanabe Nara Institute of Science and Technology (NAIST), Japan hayashi.kazuki.hl4@is.naist.jp ozaki.shintaro.ou6@naist.ac.jp {sakai.yusuke.sr9, kamigaito.h, taro.watanabe}@is.naist.jp
1 Introduction
Large-scale Vision-Language Models (LVLMs) liu2024llavanext; abdin2024phi3technicalreporthighly; ye2024mplugowl3longimagesequenceunderstanding; bai2025qwen25vltechnicalreport extend language models with visual perception and enable unified multimodal reasoning. As LVLMs move from research prototypes to real-world interactive systems, they are increasingly deployed in navigation, education, and assistive technologies pmlr-v229-zitkovich23a; jimaging10050103; morita2024contextawaresupportcolorvision; zhou2025autovla. In such applications, LVLMs serve as a medium through which users access and interpret visual information in their environment.
Figure 1: An example of communication failure due to CVDs: a user with CVD may not correctly interpret the model’s color-based instructions. This highlights the need for LVLMs to account for perceptual differences in real-world interactions.
A key yet underexplored assumption in such deployments is that users share a similar visual experience of the world. In reality, visual perception varies across individuals. Color vision in particular shows substantial variability due to biological factors Emery2019-hi, and Color Vision Deficiencies (CVDs) are a common form of such diversity, impairing discrimination between colors such as red and green Birch2012-xg; roberson2007color. These differences are not only perceptual but also functional: color vision affects wayfinding, diagram and chart reading, medical interpretation, and user interface navigation. As illustrated in Figure 1, an LVLM-based navigation assistant instructing a traveler to “Take the green line” may fail to provide actionable information for a red–green color blind user, raising concerns about inclusion, accessibility, and fairness in real deployments Kawakita2024. Systems designed around a “standard” visual experience may silently exclude users with different perceptual characteristics, even when linguistic content remains fully accessible.
Current LVLMs are trained largely on web-scale image-text corpora that implicitly assume typical color vision paik-etal-2021-world; samin-etal-2025-colorfoil; Rahmanzadehgervi_2024_ACCV. As a result, these models acquire substantial knowledge about color, color blindness, and clinical diagnostic tools. However, whether such knowledge enables the model to simulate alternative visual percepts remains unclear. This distinction is central: explaining what CVD is in language is not the same as perceiving colors as a CVD user would. To our knowledge, the ability of LVLMs to reproduce perceptual diversity has not been systematically investigated.
To address this, we leverage the Ishihara Test ishihara1917test, a clinically validated tool widely used for diagnosing CVDs in humans. Rather than using Ishihara solely as a screening instrument, we repurpose it as a diagnostic probe for LVLMs. Because Ishihara plates are designed to elicit different perceptual responses across color vision types, they allow direct comparison between model outputs and established human perceptual categories. In addition, the test provides a standardized visual stimulus that isolates color-based perceptual variation without cultural or linguistic interference, and its clinical validation over a century allows perceptual differences to be evaluated without new annotation pipelines or subjective rating procedures.
We evaluate LVLM behavior through three complementary perspectives. At the generation level, we assess whether the model outputs the digit perceived by users with different CVD types. At the confidence level, we analyze perplexity under simulated CVD conditions to determine whether uncertainty patterns align with human perceptual tendencies. At the representation level, we apply a layer-wise LogitLens probing nostalgebraist2020logitlens method to assess whether internal activations encode distinctions between vision types. Together, these perspectives allow us to analyze how perceptual variation interacts with LVLMs’ generative, probabilistic, and latent processing mechanisms.
Our findings show that although LVLMs possess strong linguistic knowledge about CVDs and can describe the Ishihara Test, they fail to reproduce the perceptual outcomes experienced by individuals with CVD. Even when given additional linguistic descriptions or visual exemplars, the models default to normative color perception, indicating insufficient alignment between linguistic knowledge and perceptual grounding. These results suggest that current LVLMs lack mechanisms for representing alternative perceptual experiences, which has implications for accessibility, fairness, and human-centered deployment in multimodal AI as such systems increasingly mediate visual information.
2 Background
2.1 Color Vision Deficiency and Ishihara Test
Color vision in humans relies on three types of cone photoreceptors: red, green, and blue, each sensitive to long, medium, or short wavelengths. CVDs affect about 8% of males and 0.5% of females Birch2012-xg. When one type of cone is missing or doesn’t work properly, large parts of the color spectrum lose their typical hue or brightness, resulting in CVDs. The three main types are summarized below:
Protanopia
Absence of long wavelength (L) cones makes reds appear very dark and compresses the differences among many reds, greens, and browns; purples are often mistaken for blue.
Deuteranopia
Absence of medium wavelength (M) cones leaves brightness almost unaffected but shifts greens toward beige, producing strong red and green confusion.
Tritanopia
Absence of short wavelength (S) cones narrows the blue to yellow axis; blues drift toward green, cyans toward gray, and yellows toward pink or light gray, greatly reducing blue versus yellow separability.
Plate Type Answer
[Uncaptioned image] Normal 12
Protanopia 12
Deuteranopia 12
Tritanopia 12
[Uncaptioned image] Normal 8
Protanopia 3
Deuteranopia 3
Tritanopia 8
[Uncaptioned image] Normal 26
Protanopia 6
Deuteranopia 2
Tritanopia 26
Table 1: Ground-truth answers for each Ishihara plate under different vision conditions, based on clinical interpretations of the Ishihara color test ishihara1917test.
To screen for CVDs, the Ishihara Test (ishihara1917test) has been widely used for decades in clinical and research settings Fanlo_Zarazaga2019-wc. Each plate shows a figure made of colored dots. People with normal vision usually identify it, while those with CVDs may see a different number or nothing. For instance, the plate in Table 1 looks like “8” to a person with normal vision, but is often read as “3” by people with red and green deficiencies Birch2012-xg. By using typical color confusions across CVDs, the Ishihara Test remains a fast and reliable tool for identifying these conditions.
2.2 CVDs and Ishihara Tests in Recent Research
Despite its limited size, the Ishihara Test remains a clinically validated and reliable tool for CVD detection. CHEN1995463 applied automated analysis to recognize Ishihara patterns, enabling automated color vision testing. Later work confirmed that digital versions retain reliability across devices zhao2019comparison; klinke2024comparative. computers14020034; grover2025huemanityprobingfinegrainedvisual introduced synthetic Ishihara-style datasets to assess model performance on digit or character recognition under red and green confusion. CVDs compensation methods also use Ishihara principles to enhance accessibility for people with color vision deficiency. Key approaches include physiologically based daltonization 5290741, eigenvector-based recoloring s19102250, and neural real-time Ishihara recoloring inbook. These works highlight Ishihara plates as diagnostic tools and foundations for assistive color algorithms.
Input Extracted mPLUG-Owl3 Output
[…]
Table 2: Excerpts from mPLUG-Owl3 showing knowledge of the Ishihara test and CVDs. The full outputs are available in Appendix B.
2.3 Color Understanding in LVLMs
Early VQA models balanced_vqa_v2 explored color-related questions but relied on priors over visual understanding, yielding inconsistent results. Agrawal_2018_CVPR; Kervadec_2021_CVPR proposed debiased datasets to reduce language bias and improve visual grounding. Transformer-based models pmlr-v139-radford21a; pmlr-v139-kim21k enhanced image-text alignment and performed well on some tasks yet still struggle with fine-grained color distinctions samin-etal-2025-colorfoil. Although recent LVLMs exhibit strong visual reasoning, their color understanding remains limited by insufficient supervision for subtle or ambiguous cases. burapacheep2024colorswapcolorwordorder; shahgir2024illusionvqachallengingopticalillusion found that models struggle with basic color-word grounding and are easily misled by perceptual illusions. liang2025colorbenchvlmsunderstandcolorful evaluated diverse color tasks and showed even large models struggle with shifts and ambiguity, revealing persistent limitations.
3 Experimental Setup
3.1 Pilot Study
Before analyzing our main question of whether LVLMs can simulate aspects of human color perception, we first verified that they possess basic language-level knowledge about color vision deficiencies and the Ishihara Test. To test this prerequisite, we used two types of prompts: (i) asking an Ishihara plate “What is this image used for?”, and (ii) requesting definitions such as “What is Protanopia as a type of color vision deficiency?”. As shown in Table 2, all evaluated models were judged to respond correctly. This confirms that they satisfy the necessary prerequisite knowledge for the analyses that follow. Appendix B provides full outputs.
3.2 Research Questions
We use the Ishihara Test as with humans to analyze how LVLMs simulate color vision across task settings [^1] We use 25 out of 38 plates that require numeral responses to ensure consistent analysis across visual conditions.. In our research, “color perception” means the model’s ability to respond like humans with specific CVDs, rather than identifying colors. We investigate this across three levels: generation, confidence, and internal representation.
Type Prompt
Base {Image: Ishihara plate}
You are Protanopic.
What number do you see? Only provide the answer. No additional information.
Linguistic Support {Image: Ishihara plate}
You are Protanopic.
Protanopia means you have difficulty distinguishing red tones, as red appears dimmer and may be confused with green or brown.
What number do you see? Only provide the answer. No additional information.
Visual Support {Reference Image: Ishihara plate}
You are Protanopic.
Example 1: Image shows an Ishihara plate. A Protanopic person sees this number: 2.
{Image: Ishihara plate}
Now, look at the following image. What number do you see? Only provide the answer. No additional information.
Doctor-Style {Image: Ishihara plate}
[…]
Choose from the following options:
- Normal - Protanopia - Deuteranopia - Tritanopia Please output your answer.
Table 3: Examples of prompts used in our experiments. Base, Linguistic, and Visual prompts are used in RQ1-RQ2 for digit generation, and Doctor-Style is used in RQ3 for vision-type inference. All include an Ishihara plate as visual input.
RQ1 (Generation level).
[…]
RQ2 (Confidence level).
Do model confidence patterns vary across vision conditions, and do uncertainty signals (measured via perplexity) reflect condition-specific perceptual ambiguity observed in humans? This probes whether LVLMs assign different likelihoods to Ishihara responses depending on simulated vision type.
RQ3 (Representation level).
Do internal layerwise activation patterns encode distinctions consistent with color vision deficiency categories? This examines whether alternative perceptual states are represented at the latent level, even if not expressed in the final output.
3.3 Task Definition
The task is inspired by the human Ishihara Test: given an Ishihara plate, the model is instructed to simulate a target vision condition and report the perceived digit. We simulate four conditions: {Normal, Protanopia, Deuteranopia, Tritanopia}. Since the Ishihara Test does not assess Tritanopia, the correct digits for Normal and Tritanopia coincide. We exploit this property by treating Tritanopia as a control condition: if a model meaningfully incorporates the instruction “You are Tritanopic,” its output distribution or confidence should differ from Normal despite identical ground-truth digits. As shown later, most LVLMs instead return nearly identical responses for the two conditions, suggesting weak sensitivity to instructed perceptual states.
Prompts
To analyze whether LVLMs can simulate CVDs under different instructional prompts, we use Ishihara plates and design four types of prompts: Base, Linguistic Support, Visual Support and Doctor-Style, as shown in Table 3. We tested multiple variants within each prompt type but observed no substantial differences. Thus, we use one representative set in our main experiments and report the variant analysis in Appendix C.
Base Prompts
These provide only the condition, e.g., “You are Protanopic”, and ask the model to report the number in the Ishihara plate. This tests whether the model can simulate CVD perception specified by the condition.
Linguistic Support
In addition to the Base prompt, these add a short description of the impairment, e.g., “red tones appear darker,” to assess whether textual context aids prediction.
Visual Support
These include a brief condition description and a reference example showing what number appears in another plate, before asking the model to identify the number in a new image. This tests whether few-shot visual examples help simulate CVD perception.
Doctor-Style
Instead of generating digits, the model infers the vision type (Normal, Protanopia, Deuteranopia, or Tritanopia) from a given response. Used in RQ3, this prompt avoids tokenization issues with multi-digit numerals (e.g., “18” $\rightarrow$ “1”, “8”) by using diagnosis labels, which are consistently tokenized and clearly distinguishable across models.
Normal Protanopia Deuteranopia Tritanopia Avg.
LVLM Size Base Ling. Vis. Base Ling. Vis. Base Ling. Vis. Base Ling. Vis. Base Ling. Vis.
Llama-3.2 11B 14.3 9.5 5.9 5.9 11.8 5.9 4.8 23.8 9.2 11.3
LLaVA-NeXT 13B 52.4 52.4 5.9 5.9 5.9 5.9 52.4 61.9 29.2 31.5
mPLUG-Owl3 7B 71.4 71.4 62.0 5.9 5.9 5.9 5.9 5.9 17.7 66.7 66.7 62.0 37.5 37.5 36.9
Phi-3.5 4.2B 0.0 0.0 4.8 0.0 0.0 5.9 0.0 0.0 11.8 0.0 0.0 9.5 0.0 0.0 8.0
Qwen2.5-VL 7B 28.6 0.0 4.8 0.0 0.0 5.9 0.0 0.0 0.0 23.8 0.0 4.8 13.1 0.0 3.9
GPT-4o 90.5 90.5 90.5 23.6 17.7 23.6 5.9 23.5 5.9 71.4 66.7 42.9 47.9 48.4 40.7
Table 4: Digit Accuracy (%) of each LVLM under simulated Vision types: Normal and three CVD types, with support settings: Base, Ling. (Linguistic), and Vis. (Visual). Vis. was used only for models supporting multi-image input. Plates for which no correct digit exists (marked as “N/A”in the Gold answers in Table 5) were excluded from the accuracy calculation.
3.4 Evaluation Metrics
We quantify each research question using a dedicated and interpretable metric for each aspect.
Digit Accuracy (RQ1)
The proportion of Ishihara plates on which the model’s generated digit matches the ground‑truth numeral under the specified vision condition.
Per‑Token Perplexity (RQ2)
For each plate, we concatenate the image, prompt, and gold numeral answer into a single input sequence. All tokens preceding the answer are masked with ignore_index, ensuring that only the answer tokens contribute to the loss. The answer consists of $L$ tokens, denoted as $\boldsymbol{x}_{1:L}$. We compute their token‑level log‑probabilities
$v_{i}=\log P\left(x_{i}\mid\textit{prompt},\,\boldsymbol{x}_{1:i-1}\right),$
and the average negative log‑likelihood
$\mathcal{L}\;=\;-\frac{1}{L}\sum_{i=1}^{L}v_{i},\quad\mathrm{PPL}\;=\;\exp(\mathcal{L}).$
A lower $\mathrm{PPL}$ indicates higher confidence in the forced-decoded answer.
Layer‑Wise Diagnosis Probability (RQ3)
This approach builds on prior work showing that intermediate representations influence model predictions wendler-etal-2024-llamas; schut2025multilingual. At each Transformer layer, we project the hidden state $\boldsymbol{h}$ into the vocabulary space using the unembedding matrix $\mathbf{W}^{\text{unembed}}$ and apply a softmax:
$\mathrm{LogitLens}(\boldsymbol{h})=\mathrm{softmax}\left(\mathbf{W}^{\text{unembed}}\,\boldsymbol{h}\right).$
The resulting distribution gives the probability of every token at that layer nostalgebraist2020logitlens. For a specific diagnosis‑label token $t$ (e.g., Protanopia or Deuteranopia) we record
$P(t)=\frac{\exp\left(\boldsymbol{W}_{t}^{\text{unembed}}\cdot\boldsymbol{h}\right)}{\sum_{j}\exp\left(\boldsymbol{W}_{j}^{\text{unembed}}\cdot\boldsymbol{h}\right)}.$
Plotting $P(t)$ across layers reveals how strongly the model favors each color‑vision diagnosis while processing the plate, offering insights into its internal decision-making process.
3.5 Models
We evaluate six LVLMs: Phi-3.5 abdin2024phi3technicalreporthighly, Qwen2.5-VL bai2025qwen25vltechnicalreport, mPLUG-Owl3 ye2024mplugowl3longimagesequenceunderstanding, LLaVA-NeXT liu2024llavanext, Llama-3.2 grattafiori2024llama3herdmodels, and GPT-4o openai2024gpt4ocard. GPT-4o is proprietary; the remaining five are open-source and were selected for passing the Pilot Study prerequisite. Appendix A provides the detailed settings.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
Llama 19 8 8 42 7 1 2 4 15 2 42 42 1 7 7 4 53 15 15 3 4 52 52 52 8
LLaVA 12 8 6 29 37 3 6 16 7 2 6 9 5 6 7 16 3 1 1 1 1 28 42 35 96
mPLUG 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 100 100 10 100 23 42 35 96
Phi
Qwen 12 8 6 29 57 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 48 74 95
GPT 12 8 6 29 57 5 3 15 74 2 6 74 45 5 7 16 73 74 74 74 74 5 42 35 96
Protanopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 6 2 5 6
Llama 19 2 5 17 3 3 2 6 3 2 2 2 45 7 4 6 46 73 4 53 52 4 4 4 2
LLaVA 12 8 6 29 57 3 6 16 7 2 6 9 1 6 7 16 3 1 1 1 1 20 42 38 96
mPLUG 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 1 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 74 7 74 74 74 74 38 43 74 95
GPT 12 3 3 29 55 3 5 15 4 9 3 37 45 N/A 7 16 13 74 2 74 27 25 N/A 39 96
Deuteranopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 2 4 3 9
Llama 19 2 5 52 1 53 52 49 45 53 52 52 36 55 7 56 4 4 2 49 53 42 52 37 8
LLaVA 12 8 6 29 57 3 3 16 14 2 6 9 10 6 7 16 3 10 1 1 12 20 42 38 96
mPLUG 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 1 1 10 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 53 7 74 74 74 74 88 43 30 95
GPT 12 8 6 29 57 3 8 N/A 74 2 6 97 45 3 7 16 73 74 N/A N/A 5 93 42 35 96
Tritanopia Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
Llama 19 2 5 17 1 3 2 4 1 3 4 42 46 4 7 3 8 4 4 4 1 3 4 3 8
LLaVA 12 8 6 29 37 3 3 16 7 2 6 9 1 6 7 16 3 1 1 1 12 28 42 38 96
mPLUG 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 10 10 10 10 23 42 35 96
Phi
Qwen 21 8 6 29 37 74 74 74 74 74 74 74 74 5 3 53 3 74 0 74 74 38 43 30 95
GPT 12 8 6 29 57 N/A 3 15 74 2 3 37 36 5 7 16 33 74 74 6 3 N/A 42 35 96
Table 5: Predicted outputs of six LVLMs on Ishihara color-vision test plates under four simulated conditions (Normal, Protanopia, Deuteranopia, and Tritanopia) using the Base prompt only. The “Gold” row shows the ground-truth digits. Plates marked as “N/A” indicate those without a correct digit (e.g., blank or pattern-only plates) and are excluded from accuracy calculations.
4 Results and Analysis
4.1 RQ1: Generation Level
We use Digit Accuracy to evaluate how often the model correctly identifies the numeral on each Ishihara plate. Table 4 shows accuracy across conditions and prompt types. In the Normal condition, humans are expected to score 100 %, but only GPT-4o approaches this with 90.5 %, followed by mPLUG-Owl3 at 71 %, while some models score near 0 %. Performance collapses for both Protanopia and Deuteranopia, with no model exceeding 24 %, clearly indicating the difficulty of simulating CVDs. With Linguistic Support, brief descriptions shift accuracy by at most two points without changing rankings; GPT-4o improves to 23 % on Deuteranopia but drops on Tritanopia. Qwen2.5-VL often refuses under disability-related terminology, averaging 0 %. With Visual Support, few-shot examples have minimal effect; accuracy shows no consistent improvement and sometimes decreases.
Table 5 lists the gold digits and Base-prompt predictions across 25 plates. GPT-4o reproduces most digits under Normal vision, and mPLUG-Owl3 performs well, but the remaining models are inconsistent. Under Protanopia and Deuteranopia, most models repeat Normal outputs, return empty strings, or collapse to a single digit, e.g., “70.” Thus, strong Normal accuracy does not imply successful CVD simulation. Even with Linguistic or Visual Support, accuracy improves little and sometimes introduces new errors, indicating that prompt modifications alone cannot reproduce perceptual variation. Tritanopia shares the same ground-truth digits as Normal and therefore serves as a control condition. Models output nearly identical digits for both, matching clinical answers but indicating fallback to Normal rather than simulation. Severe failures under Protanopia and Deuteranopia further show that agreement in Tritanopia does not imply perceptual simulation.
At the generation level, current LVLMs assume normal color vision and, even with prompt support, fail to generate text that matches CVDs perception.
LVLM Size Prompt Normal Protanopia Deuteranopia Tritanopia
Mean SD Mean SD Mean SD Mean SD
mPLUG‑Owl3 7B Base 3.8 6.7 69.7 78.8 69.0 116.0 3.7 6.3
Linguistic 3.8 6.7 78.9 101.5 105.9 185.6 4.2 6.9
Visual 16.0 40.2 36.5 43.3 37.1 59.4 14.3 33.5
Phi‑3.5 4.2B Base 14.5 13.3 140.5 325.5 80.0 256.1 23.9 28.6
Linguistic 14.5 13.3 334.0 792.7 158.2 544.1 24.9 28.7
Visual 7.0 6.2 13.1 12.3 12.7 12.8 7.6 7.3
Qwen2.5‑VL 7B Base 46.0 83.3 72.8 62.1 57.2 69.1 27.4 36.6
Linguistic 46.0 83.3 62.0 68.7 51.4 59.7 27.7 36.6
Visual 45.0 56.5 313.7 381.8 131.7 172.0 78.0 82.1
Table 6: Mean and standard deviation (SD) of per-token perplexity across four simulated vision conditions (Normal, Protanopia, Deuteranopia, Tritanopia) under three prompt settings (Base, Linguistic, Visual) for each LVLM. Only models that could be tested under all three settings are shown.
Figure 2: Violin plots of mPLUG-Owl3 perplexity on Ishihara digit prediction. Columns indicate the condition specified in the prompt (Normal, Protanopia, Deuteranopia, Tritanopia). Within each column, colors denote the vision type that defines the candidate digit whose perplexity is measured (Normal: blue, Protanopia: red, Deuteranopia: green, Tritanopia: yellow). Lower perplexity (for the gold answer) indicates higher model confidence.
4.2 RQ2: Confidence Level
To evaluate model confidence, we measure per-token perplexity for each forced-decoded output. Comparing perplexity across conditions reveals whether confidence aligns with accuracy, whether linguistic or visual cues reduce uncertainty, and how strongly models default to normal vision.
Table 6 reports the mean and standard deviation (SD) of perplexity (shown as mean $\pm$ SD) for all models under the three prompt settings. Across the board, Normal and Tritanopia remain lower than Protanopia and Deuteranopia, indicating a pronounced bias toward normal vision, though exceptions are observed in Llama-3.2. One model rises from about $4\pm 7$ on Normal to $70\pm 80$ on Protanopia, while another climbs from about $14\pm 13$ to $140\pm 330$. Perplexity generally increases under Linguistic and Visual Support. Although visual cues lower perplexity for a few models, the tendency persists: Normal is always lowest and the red and green deficiencies remain highest.
LVLM Size Normal Protanopia
Llama-3.2 11B 36.0 72.4
LLaVA-NeXT 13B 20.0 86.2
mPLUG-Owl3 7B 28.0 86.2
Phi-3.5 4.2B 100.0 3.5
Qwen2.5-VL 7B 4.0 86.2
Table 7: Doctor-style diagnosis accuracy (%) under Normal and Protanopia conditions for various LVLMs.
Figure 3: Layer-wise diagnosis probabilities for LLaVA, averaged over all plates.
Normal Protanopia Deuteranopia Tritanopia Avg.
LVLM Size Base Ling. Vis. Base Ling. Vis. Base Ling. Vis. Base Ling. Vis. Base Ling. Vis.
mPLUG-Owl3 7B 71.4 71.4 62.0 5.9 5.9 5.9 5.9 5.9 17.7 66.7 66.7 62.0 37.5 37.5 36.9
Qwen2.5-VL 7B 28.6 0.0 4.8 0.0 0.0 5.9 0.0 0.0 0.0 23.8 0.0 4.8 13.1 0.0 3.9
GPT-4o 90.5 90.5 90.5 23.6 17.7 23.6 5.9 23.5 5.9 71.4 61.9 42.9 47.9 48.4 40.7
MedVLM-R1 57.1 57.1 0.0 5.9 5.9 0.0 5.9 5.9 0.0 61.9 61.9 0.0 32.7 32.7 0.0
Qwen‑DotPattern 71.4 71.4 23.8 5.9 5.9 5.9 5.9 5.9 17.7 66.7 66.6 23.8 37.5 37.5 17.8
Table 8: Digit accuracy (%) of each LVLM under simulated vision types: Normal, Protanopia, Deuteranopia, and Tritanopia, with support settings: Base, Ling., and Vis. As for Vis., it was used only for models supporting multi-image input. This table extends Table 4 by including additional results for MedVLM-R1 and Qwen-DotPattern.
Figure 2 plots perplexity for mPLUG-Owl3. We selected this model because it supports multiple images and exhibits the most stable perplexity across prompt settings. The three prompt settings share nearly identical shapes; Normal clusters near zero, whereas Protanopia and Deuteranopia rise to tall peaks. Adding a brief description or a reference image leaves the distribution almost unchanged, which shows that prompt tweaks do not alter the model’s internal uncertainty. Other results are in Appendix E, showing the same trend: low perplexity for Normal vision, higher and variable for CVDs, with little effect from prompt changes.
At the confidence level, Normal vision consistently yields the lowest perplexity, while the CVD conditions result in similarly high perplexity, suggesting that models are fundamentally confused when simulating altered color vision and fail to reproduce such perceptual states.
4.3 RQ3: Internal Representation Level
We analyze whether LVLMs can perform doctor-style diagnosis by inferring vision types from responses using the Doctor-Style prompt (Table 3). This setting evaluates whether models can identify vision types from a response, analogous to how a doctor interprets Ishihara results. For each plate, we force-decode four labels (Normal, Protanopia, Deuteranopia, and Tritanopia) and select the most probable one. As shown in Table 7, models perform well on Protanopia (over 70%, except Phi-3.5), but poorly on Normal (under 30%). This contrasts with Table 4, where they failed to recognize digits under red and green conditions.
LVLM Size Prompt Normal Protanopia Deuteranopia Tritanopia
Mean SD Mean SD Mean SD Mean SD
Qwen2.5‑VL 7B Base 46.0 83.3 72.8 62.1 57.2 69.1 27.4 36.6
Linguistic 46.0 83.3 62.0 68.7 51.4 59.7 27.7 36.6
Visual 45.0 56.5 313.7 381.8 131.7 172.0 78.0 82.1
MedVLM-R1 2B Base 15.5 28.6 166.4 325.3 86.1 160.8 14.7 26.2
Linguistic 15.5 28.6 189.7 369.6 91.5 195.2 13.2 24.5
Visual 48.6 63.6 83.7 92.3 72.7 89.6 50.9 62.1
Qwen‑DotPattern 2B Base 4.7 8.9 79.2 136.3 75.4 171.7 4.7 6.6
Linguistic 4.7 8.9 106.3 213.0 100.2 249.9 4.7 6.5
Visual 6.2 9.1 41.5 55.2 28.7 47.1 7.4 9.9
Table 9: Mean and standard deviation (SD) of per-token perplexity across four simulated vision conditions under three prompt settings (Base, Linguistic, Visual). This table extends Table 6 by including additional results for MedVLM-R1 and Qwen-DotPattern.
LVLM Size Normal Protanopia
Qwen2.5-VL 7B 4.0 86.2
MedVLM-R1 7B 21.4 88.0
Qwen-DotPattern 7B 85.7 0.0
Table 10: Doctor-style diagnosis accuracy (%) under Normal and Protanopia conditions for MedVLM-R1 and Qwen-DotPattern, extending Table 7.
To further examine this behavior, we use Layer-Wise Diagnosis Probability, inspired by Logit Lens nostalgebraist2020logitlens, to track token-level probabilities across transformer layers. Figure 3 uses LLaVA-NeXT as a representative example. Results for other models are provided in Appendix F, where no consistent trends were observed across models, indicating that the layer-wise behavior varies depending on the architecture. LLaVA-NeXT tends to interpret images as Normal in mid-layers, shifting to Protanopia at the final layer, suggesting inconsistent processing. Nearly identical curves across ground-truths imply poor differentiation between conditions.
At the internal representation level, the model exhibits nearly identical behavior across all color vision types, showing no evidence of reasoning that reflects differences between conditions. This indicates a fundamental inability to understand and reproduce variations in color perception.
4.4 Domain Knowledge and Pattern Memorization
To deepen our analysis, we conducted two focused experiments as outlined below.
Effect of medical expertise
As Ishihara plates originate in clinical practice, we expect that an LVLM trained on medical images and reports would outperform standard models. We therefore evaluated MedVLM‑R1 pan2025medvlmr1incentivizingmedicalreasoning, which is built on Qwen‑2.0‑VL wang2024qwen2 and fine‑tuned on medical data, to investigate whether training on clinical data improves performance on color-based diagnostic tools.
Dot pattern memorization hypothesis
A potential concern is that models may solve the task by exploiting superficial dot-pattern regularities in Ishihara plates, rather than demonstrating perceptual reasoning. To directly probe this, we fine‑tuned Qwen2.0‑VL on 10,000 synthetic Ishihara‑style images released by grover2025huemanityprobingfinegrainedvisual, producing a model specialized in dot‑pattern recognition. We refer to this model as Qwen‑DotPattern. This setup allows us to assess whether pattern memorization alone can explain model performance. Training details are provided in Appendix G.
Table 8 shows generation-level accuracy for MedVLM-R1 and Qwen-DotPattern. Under the Normal condition with the Base prompt, both outperform the Qwen2.5-VL baseline (MedVLM-R1: 57.1%, Qwen-DotPattern: 71.4%), indicating that domain-specific training with dot-pattern stimuli enhances recognition under typical vision. However, under Protanopia and Deuteranopia, follow the same trend as other models, showing low accuracy across prompts, suggesting that neither medical knowledge nor pattern memorization reproduces CVD perceptual distortions.
Table 9 presents the confidence level results in terms of perplexity. Under the Normal condition, both MedVLM-R1 and Qwen-DotPattern show lower perplexity than Qwen2.5-VL, indicating that domain adaptation enhances confidence on typical inputs. However, under red–green CVDs, both show higher perplexity than the base model, suggesting that medical or synthetic training improves confidence in normal vision but increases miscalibration when color perception is impaired.
We extend the doctor diagnosis analysis by adding MedVLM-R1 and Qwen-DotPattern, as shown in Table 10. MedVLM-R1 shows similar accuracy in both Normal (21.4%) and Protanopia (88.0%) settings, consistent with general model trends. This suggests that while medical training supports classification under impaired vision, it does not improve Normal vision. Qwen-DotPattern shows the opposite pattern. It performs well under Normal (85.7%) but fails under Protanopia (0.0 %). This suggests the model may rely on memorized dot layouts rather than learning generalizable structure, improving recognition in standard settings but reducing adaptability under altered vision.
These results show that training on domain-specific and dot-pattern images can improve model behavior in normal color vision, but do not lead to gains in reproducing the perceptual experience of CVDs. This highlights the value of the Ishihara Test, which remains a strong diagnostic benchmark despite its small size, as it is unaffected by pattern memorization and reliably evaluates a model’s reasoning under impaired color perception.
5 Conclusion
This work examined whether LVLMs can simulate variation in human color perception using the Ishihara Test. We analyzed model behavior at the levels of generation, confidence, and internal representation to assess whether linguistic knowledge about CVDs reflects perceptual grounding. Although LVLMs can describe CVDs and the test procedure, they default to a normative color percept and do not reproduce the digit responses associated with different vision types. Additional textual guidance or visual manipulation does not induce alternative percepts, and probing reveals only weak separation between vision types in internal activations. These findings indicate that current LVLMs lack mechanisms for alternative perceptual experiences, highlighting the need for approaches that support perceptual diversity in future multimodal systems for accessible and human-centered deployment.
6 Limitations
Dataset Scope and Scale.
Our evaluation relies on a small set of Ishihara plates that primarily target red–green color-vision deficiencies. The limited number and narrow focus reduce the diversity of conditions tested, so our findings may not generalise to other CVD types or broader visual tasks. While scaling to broader datasets and real-world tasks remains important future work, our focus here is on establishing a methodology and analysing behavioural differences across prompted vision conditions rather than memorisation. Possible pre-training exposure to Ishihara plates would mainly affect the Normal condition and does not account for failures under CVD prompts.
Limitations in Evaluating Tritanopia.
The Ishihara Test is primarily designed to detect red–green deficiencies (Protanopia and Deuteranopia) and is not suitable for assessing blue–yellow deficiencies such as Tritanopia. Moreover, the medically defined ground-truth answers for Tritanopia coincide with those for Normal vision. In our setting, Tritanopia therefore functions largely as a control for instruction following rather than as a genuine benchmark for blue–yellow deficiencies. As a result, conclusions on Tritanopia are limited, and future work should evaluate all major CVD types using more general tasks beyond Ishihara plates.
Evaluation Methodology.
We prompt LVLMs to simulate color-blind perception in a controlled diagnostic setting, tasking them with identifying digits in Ishihara plates. We use a restricted set of prompts (role-playing, doctor-perspective, step-by-step). While failures were broadly consistent across them, the prompt space is incomplete and other formulations may elicit different behaviours. Consequently, our conclusions characterise LVLM behaviour under this evaluation format rather than ruling out prompts that could induce more faithful CVD simulation. We also treat medically defined Ishihara digits as ground truth without validation from individuals with CVD or clinicians, which may limit ecological validity.
Generalisability of Model Behavior.
Performance on stylised Ishihara plates does not imply that a model can handle color perception in broader settings. Models may rely on dot-pattern heuristics or memorised cues rather than simulating color confusion. Prior work also shows that vision–language models can exhibit color biases raj-etal-2024-biasdora, suggesting that controlled diagnostic performance may not translate to broader color understanding. Caution is therefore required when extrapolating to natural images. Finally, this study is diagnostic rather than prescriptive: we do not test mitigation strategies, and developing such methods remains an important direction for CVD-aware LVLMs.
7 Ethical Considerations
Inclusive Design for Perceptual Diversity.
Recent AI research communities, including ACL, have emphasized the importance of designing fair and inclusive systems that accommodate users with perceptual or cognitive differences. Color vision deficiency, which affects a significant portion of the global population, is one such factor that is often overlooked in model development and evaluation. By explicitly testing LVLM behavior under CVD conditions, this study contributes to a broader understanding of accessibility in multimodal AI systems and encourages further research in this direction. Building on this motivation, we also discuss fairness-related concerns in the context of model evaluation.
Not for Diagnostic Use.
This study aims to evaluate whether models can simulate aspects of color vision deficiency as a step toward developing systems that better capture the diversity of human perception. It is not intended to replace professional diagnosis or support clinical decision-making. Even if a model produces correct responses to certain vision tests, these outputs must not be used for medical purposes or institutional assessments. To prevent such misuse, we clearly define the scope and limitations of our evaluation and explicitly state that the outputs are not a substitute for expert judgment in clinical contexts.
Licensing and Copyright.
The Ishihara Test ishihara1917test is a long-established diagnostic tool widely used in academic research across fields such as ophthalmology, psychology, education, and machine learning (zhao2019comparison; klinke2024comparative; s19102250). We did not reproduce, modify, or redistribute the plates; they were used solely for research and analysis.
AI-assistant Usage.
Portions of the manuscript, such as prompt templates and wording adjustments, were drafted with the assistance of GPT-4 to ensure linguistic clarity. All technical content, analysis scripts, and final decisions were made by us.
Appendix A Detailed Model Settings
In this study, to ensure fair and consistent performance comparisons, all experiments were conducted using a single NVIDIA RTX 6000 Ada GPU. All models were run under the same computational settings, with generation performed using half quantization to reduce memory usage while maintaining model fidelity. We used publicly available models from Hugging Face and OpenAI, as summarized in Table 11. Generation was conducted using greedy decoding, with fixed random seeds to ensure reproducibility. All other settings adhered to the default configurations of each model implementation.
Model Params HuggingFace Name / OpenAI Name
Phi3.5 4.2B microsoft/Phi-3-vision-128k-instruct
Qwen2.5-VL 7B Qwen/Qwen-2.5-VL-7B-Instruct
mPLUG 7B mPLUG/mPLUG-Owl3-7B-240728
Llama3.2 11B meta-llama/Llama-3.2-11B-Vision-Instruct
LLaVA-NeXT 13B llava-hf/llava-v1.6-vicuna-13b-hf
MedVLM-R1 2B JZPeterPan/MedVLM-R1
GPT-4o gpt-4o-2024-11-20
Table 11: Detailed model names.
Appendix B Language Knowledge of the Ishihara Test and CVDs
Before testing visual ability, we verified that all LVLMs possess basic language knowledge of the Ishihara Test and the three major color vision deficiencies. As the main paper includes only one example, this appendix provides the full outputs for the remaining models.
Tables 1923 show representative answers from Llama-3.2, LLaVA-NeXT, mPLUG-Owl3, Phi-3.5, and Qwen2.5-VL. Each table has input prompts on the left and trimmed model responses on the right. All models except Phi-3.5, which returned an error for the Ishihara query, accurately describe the test and define Protanopia, Deuteranopia, and Tritanopia, confirming sufficient textual knowledge. These results show that failures in the main experiments arise from weaknesses in visual grounding, not language understanding.
Prompts in Detail
For each color vision condition, we designed three types of prompts: Base, Linguistic Support, and Visual Support. The full prompt templates are provided in Tables 1416.
Table 14 presents the Base prompts, which specify only the condition (e.g., “You are Protanopic”) and then ask the model to report the number shown in the Ishihara plate. This represents the simplest setting, testing whether the model can directly simulate the specified condition.
[…]
Table 16 shows the Visual Support prompts. In this case, the condition is accompanied by a few shot reference example (e.g., “A Protanopic person sees this number: {few_shot_number}”) before asking the model to identify the number in a new plate. This setup tests whether visual exemplars serve as useful guidance for simulating CVD perception.
Together, these tables make explicit how the three prompt types were implemented, complementing the overview provided in Section 3.3.
Appendix C Effects of Prompt Variants on Simulating CVD Perception
Prompt phrasing can influence how LVLMs interpret images, particularly when the task requires aligning visual input with an altered perceptual perspective. To assess whether failures in our main experiments were attributable to prompt design rather than perceptual grounding, we evaluated four alternative prompt formulations for the Protanopia condition:
  • Prompt 1 (Description-Based): “What number would a Protanopic viewer likely see in this image? Only provide the predicted number.”
  • Prompt 2 (Doctor Perspective): “You are an eye doctor evaluating how different color vision types perceive this test plate. What number would a Protanopic patient report? Only output the number.”
  • Prompt 3 (Role-Play Perspective): “Answer as if you are a Protanopic viewer. Respond with the number you visually perceive. Only output the number.”
  • Prompt 4 (Step-by-Step Reasoning): “Step 1: Describe how the red and green dots differ under Protanopia. Step 2: Based on Step 1, what number appears to a Protanopic viewer? Only output Step 2.”
Table 12 reports digit accuracy under Normal and three CVD conditions for Prompts 1–3. Prompt 4 was excluded due to highly unstable behavior: models frequently produced free-form explanations, failed to emit a digit token in Step 2, or ignored the requested output format entirely.
Across all prompt variants, we observe moderate variability in the Normal condition, confirming that prompt phrasing can influence how LVLMs interpret visual input when no perceptual transformation is required. However, under Protanopia, Deuteranopia, and Tritanopia, accuracy remains low across prompts and models. Neither altering task framing (doctor vs. viewer) nor introducing explicit role-playing improved condition-appropriate responses. These results indicate that prompt formulation alone is insufficient to induce LVLMs to simulate condition-specific perceptual distortions, consistent with the main conclusion that limitations arise from insufficient visual grounding rather than linguistic control alone.
Normal Protanopia Deuteranopia Tritanopia
LVLM Size P1 P2 P3 P1 P2 P3 P1 P2 P3 P1 P2 P3
Llama-3.2 11B 9.5 4.8 9.5 17.6 11.8 0.0 11.8 11.8 11.8 4.8 4.8 4.8
LLaVA-NeXT 13B 61.9 52.4 57.1 5.9 5.9 5.9 5.9 5.9 5.9 47.6 61.9 57.1
mPLUG-Owl3 7B 71.4 76.2 66.7 5.9 5.9 5.9 5.9 5.9 5.9 66.7 71.4 66.7
Phi-3.5 4.2B 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Qwen2.5-VL 7B 28.6 28.6 38.1 5.9 0.0 0.0 0.0 0.0 0.0 14.3 19.1 28.6
Table 12: Digit accuracy (%) under three prompt variants (P1–P3) across simulated vision types. Prompt 4 is omitted due to unstable output formatting and inconsistent digit emission.
Appendix D Digit Results: Detailed Analysis
Tables 17 and 18 report the raw digit outputs that each LVLM produced for Ishihara plates under four simulated conditions (Normal, Protanopia, Deuteranopia, Tritanopia). Unless stated otherwise, plates marked as “N/A” (no correct digit) are excluded from accuracy and are discussed only for error behavior.
D.1 Linguistic Support prompts (Table 17).
Adding a short textual description yields only modest and inconsistent gains. Under Normal vision, GPT remains strong and LLaVA/mPLUG perform similarly to the Base setting, while Llama continues to miss many plates. Under simulated CVDs, condition-specific digits are seldom produced; the same plates that were difficult under Base remain challenging, and response biases (e.g., repeated 74/42) persist. Hallucinated digits also appear on N/A plates, showing that brief definitions alone are insufficient to induce condition-aware perception.
D.2 Visual Support prompts (Table 18).
Providing a few shot reference image before the target plate produces mixed results. GPT maintains strong accuracy under Normal vision but shows limited and inconsistent improvements for simulated CVDs. mPLUG performs slightly better on some plates compared to Base/Linguistic settings but still fails to reliably follow the instructed condition. Phi occasionally outputs correct answers without cross-plate consistency, while Qwen remains unstable and frequently repeats the same digits. Some regressions and hallucinated digits are also observed, indicating that single visual exemplars do not reliably induce condition-appropriate perception.
D.3 Overall Findings
Across all prompt types, LVLMs display recurring error patterns: frequent mode collapse to a small set of digits (e.g., 74, 42, 3), insensitivity to the instructed CVD condition, and hallucinated outputs on plates without valid digits. These behaviors show that neither short textual descriptions nor single visual exemplars are sufficient to ground condition-specific perception. While the models can describe CVDs accurately in text, they fail to translate these instructions into consistent digit recognition on Ishihara plates. This confirms that the main limitation lies in visual grounding rather than language knowledge or prompt design.
Appendix E Additional Perplexity Analyses
In the main text, we focused on mPLUG-Owl3 (Figure 2) because it supports multi-image input and shows the most stable perplexity across prompt settings. For completeness, we report violin plots of perplexity distributions for three additional models: LLaMA, Phi, and Qwen.
Figures 46 show perplexity measured by force decoding the Gold Answer for Ishihara digit predictions under four simulated vision conditions (Normal, Protanopia, Deuteranopia, Tritanopia) and three prompt types (Base, Linguistic, Visual).
Across these models, we observe similar tendencies as discussed in the main text: Normal vision yields consistently lower perplexity, while the CVD conditions lead to higher and more variable perplexity. Adding textual descriptions (Linguistic Support) or reference images (Visual Support) does not substantially alter these distributions. This supports the conclusion that prompt modifications are insufficient to reduce model uncertainty when simulating altered color vision.
Appendix F Detailed Diagnosis of CVDs with LVLMs
Figures 710 extend the main analysis by visualizing token-level label probabilities across transformer layers for four additional LVLMs. Each panel groups plates by the correct vision condition and plots plate-wise trajectories (thin lines) together with the average curve (bold line). Left plots correspond to Normal or Tritanopia, right plots to Protanopia or Deuteranopia.
Llama-3.2 (Figure 7), For Normal/Tritanopia, the Normal label stays near zero until three layers before the head, then rises above 50%. Red and green show an analogous late surge in the Protanopia curve, reflecting a rise in uncertainty.
mPLUG-Owl3 (Figure 8), Early layers are mixed but converge rapidly. For instance, tritanopia exceeds 80% by layer 26. When the digit is unreadable, the model shifts probability toward Tritanopia, possibly explaining its low variance under Base.
Phi-3.5 (Figure 9), Curves remain diffuse and never exceed 25% even near the output. This matches the wide perplexity spread and frequent no-answer cases.
Qwen2.5-VL (Figure 10), Normal shows local peaks rather than a smooth ascent. Red and green plates present competing red and green waves that collapse to a single token at the final layer. However, the overall probabilities remain extremely low, on the order of $10^{-6}$ even beyond layer 6, indicating very weak model confidence.
Appendix G Fine-tuning and Inference Settings
Hyperparameter Value
Model Qwen/Qwen2-VL-2B-Instruct
Torch dtype bfloat16
Seed 42
Max input length 2048
Epochs 1
Learning rate 2e-4
Batch size 1
Gradient accumulation 32
Warmup ratio 0 (not used)
Gradient checkpointing True
LoRA rank ($r$) 8
LoRA alpha 16
LoRA dropout 0.05
Save steps 200
Eval strategy None
Table 13: Hyperparameters and settings used for fine-tuning. We used Transformers wolf-etal-2020-transformers, PEFT xu2023parameterefficientfinetuningmethodspretrained, and trl libraries with default configurations unless otherwise noted.
Condition Prompt
Protanopia You are Protanopic.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Deuteranopia You are Deuteranopic.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Tritanopia You are Tritanopic.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Normal Vision What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Table 14: Base prompts for each color‐vision condition.
Setting Prompt
Protanopia You are Protanopic.
Protanopia means you have difficulty distinguishing red tones, as red appears dimmer and may be confused with green or brown.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Deuteranopia You are Deuteranopic.
Deuteranopia means you have difficulty distinguishing green tones, as green appears dimmer and may be confused with red or brown.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Tritanopia You are Tritanopic.
Tritanopia means you have difficulty distinguishing blue tones, as blue appears dimmer and may be confused with green or gray.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Normal Vision What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Table 15: Prompts with linguistic support for each color‐vision condition.
Condition Prompt
Protanopia You are Protanopic.
You have difficulty distinguishing certain colors.
Example 1:
Image shows an Ishihara plate.
A Protanopic person sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Deuteranopia You are Deuteranopic.
You have difficulty distinguishing certain colors.
Example 1:
Image shows an Ishihara plate.
A Deuteranopic person sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Tritanopia You are Tritanopic.
You have difficulty distinguishing certain colors.
Example 1:
Image shows an Ishihara plate.
A Tritanopic person sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Normal Vision You have normal color vision.
Example 1:
Image shows an Ishihara plate.
A person with normal color vision sees this number: {few_shot_number}
Now, look at the following image.
What number do you see?
Only provide the answer. If you see nothing, answer ’Nothing’. No additional information.
Table 16: Prompts with visual support examples for each color‐vision condition.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
LLama 16 8 8 42 52 4 4 52 15 3 15 42 3 4 7 4 15 15 4 15 6 3 4 52 52
LLava 12 8 6 29 37 3 6 16 7 2 6 9 5 7 7 16 3 1 1 1 1 28 42 35 96
mPlug 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 100 100 10 100 23 42 35 96
Phi
Qwen
GPT 12 8 6 29 57 5 3 15 74 2 6 74 45 5 7 16 23 26 74 29 5 26 42 35 96
Protanopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 6 2 5 6
LLama 15 8 7 42 5 6 2 3 9 8 4 2 3 4 4 3 6 2 6 4 4 6 52 2 42
LLava 12 8 6 29 37 3 6 16 7 2 6 9 1 6 7 16 3 1 1 1 1 26 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 9 1 23 42 35 96
Phi
Qwen
GPT 32 3 5 26 5 8 8 13 21 5 3 37 46 2 7 16 13 3 5 5 5 23 12 30 39
Deuteranopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 2 4 3 9
LLama 8 8 6 59 3 4 6 4 66 6 8 22 49 7 7 4 4 8 7 11 41 2 42 33 44
LLava 12 8 6 29 57 3 6 16 14 2 6 9 1 6 7 16 3 1 12 12 12 26 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 9 1 23 42 35 96
Phi
Qwen
GPT 12 3 5 29 37 2 8 13 41 2 5 97 45 3 3 16 78 6 5 3 74 93 42 35 96
Tritanopia Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
LLama 12 2 8 17 8 4 3 6 2 2 2 8 7 5 7 7 6 2 4 2 4 6 4 4 8
LLava 12 8 6 29 57 3 3 16 7 2 6 9 1 6 7 16 3 1 1 12 12 26 42 38 96
mPlug 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 10 10 10 10 23 42 35 96
Phi
Qwen
GPT 12 8 6 29 57 2 8 15 41 2 3 73 35 5 7 16 73 3 5 5 3 93 42 35 96
[…]
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
mPlug 12 8 6 29 37 5 3 15 2 6 97 90 5 5 7 16 13 N/A N/A N/A N/A 26 42 35 96
Phi 12 6 29 37 5 3 15 74 2 6 97 45 5 7 16 73 1 3 3 3 25 42 3 96 90
Qwen 8 6 38 37 74 74 15 74 74 74 3 37 74 3 16 73 4 74 42 74 74 37 24 74 74
GPT 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 33 74 74 N/A 74 93 42 35 96
Protanopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 6 2 5 6
mPlug 12 8 6 29 37 5 3 15 2 2 6 9 5 5 7 13 13 1 45 73 6 2 43 3 96
Phi 12 6 5 35 3 1 17 1 3 3 4 17 3 7 7 3 1 2 45 73 3 3 3 3 9
Qwen 7 3 8 74 7 7 3 74 N/A 74 74 74 3 3 8 N/A 7 7 74 42 4 7 3 3 7
GPT 21 3 5 26 5 5 5 15 N/A 2 3 37 45 5 N/A 16 3 5 74 6 3 25 42 33 96
Deuteranopia Gold 12 3 5 70 35 2 5 17 21 N/A N/A N/A N/A N/A N/A N/A N/A 5 2 45 73 2 4 3 9
mPlug 12 8 6 29 37 5 3 15 2 2 6 9 5 5 7 13 13 1 45 73 2 2 4 35 96
Phi 12 5 35 17 7 7 45 73 3 3 9 9
Qwen 7 7 3 74 7 7 23 74 74 74 74 74 74 74 8 74 7 7 74 42 7 7 7 7 7
GPT 14 0 5 29 37 5 9 15 74 2 3 97 45 5 N/A 16 33 2 17 74 5 23 42 35 96
Tritanopia Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
mPlug 12 8 6 29 37 5 3 15 2 6 6 90 5 5 7 13 3 N/A N/A N/A 26 42 42 35 96
Phi 12 6 23 5 3 15 74 6 9 45 7 7 73 26 42 3 96 9
Qwen 3 3 3 3 4 7 15 3 4 3 3 3 3 3 1 42 N/A N/A N/A N/A 4 3 42 3 3
GPT 21 3 6 29 N/A 5 8 15 71 2 5 57 46 3 3 16 N/A N/A N/A N/A 57 26 42 35 86
[…]
Input Model Output (LLama as an example)
What is Ishihara Test? This image is used to test color vision. It is called an Ishihara plate, named after its creator, Dr. Shinobu Ishihara, a Japanese professor of ophthalmology. The plate is a widely used tool for diagnosing color blindness, particularly red-green color blindness, which is the most common form of color vision deficiency. The plate consists of a series of dots in different colors and sizes, arranged in a specific pattern. The viewer is asked to identify a number or shape hidden within the dots. For people with normal color vision, the correct answer is obvious, but for those with color vision deficiency, the correct answer may not be apparent. In this image, the number 8̈ïs hidden among the dots. The dots are arranged in a way that creates a red 8̈s̈hape, but the dots are also scattered randomly throughout the plate, making it difficult to see the number without looking closely. For someone with normal color vision, the red 8̈s̈tands out clearly, but for someone with red-green color blindness, the number may not be visible at all. The Ishihara plate is used to diagnose color blindness because it exploits the fact that people with certain types of color vision deficiency have difficulty distinguishing between certain colors. In this case, the plate uses the fact that people with red-green color blindness have trouble distinguishing between red and green colors. By using a combination of red and green dots, the plate can help identify individuals who have difficulty seeing these colors. Overall, the Ishihara plate is a simple yet effective tool for diagnosing color blindness, and it has been widely used in medical settings for many years.
What is Protanopia? Protanopia, also known as red-blindness, is a type of color vision deficiency that affects approximately 1% of males and 0.05% of females. It is caused by a mutation in the OPN1LW gene, which codes for the long-wavelength cone cells in the retina. These cone cells are responsible for detecting long-wavelength light, which is associated with the color red. In individuals with protanopia, the long-wavelength cone cells are either absent or non-functional, leading to a reduced ability to detect red light. As a result, people with protanopia may have difficulty distinguishing between red
What is Deuteranopia? Deuteranopia is a type of color vision deficiency that affects approximately 5% of males and 0.4% of females in the United States. Deuteranopia, also known as red-green color blindness, is a type of color vision deficiency that affects a person’s ability to see certain colors, particularly red and green. People with deuteranopia have difficulty distinguishing between these two colors and may see them as similar shades of yellow or brown. Deuteranopia is caused by a deficiency in the long-wavelength cones in the retina, which are responsible for detecting red light. As a result, people with deuter
What is Tritanopia? Tritanopia is a rare form of color vision deficiency that affects a person’s ability to see the color blue. People with tritanopia have difficulty distinguishing between different shades of blue and green, and may see these colors as more yellowish or brownish. Tritanopia is a congenital condition, meaning it is present at birth, and is usually inherited in an X-linked recessive pattern. This means that the gene responsible for tritanopia is located on the X chromosome, and males are more likely to be affected because they have only one X chromosome. Females, on the other hand, are less likely to be affected
Table 19: An example of Llama’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( LLava as an example)
What is Ishihara Test? This image appears to be a color vision test, also known as an Ishihara test. The test is used to check for the presence of a condition called central serous chorioretinopathy (CSCR), which can cause central serous chorioretinopathy. The test uses a pattern of colored dots, and the presence of certain colors can indicate the presence of the condition. The pattern is typically circular and contains a number of colored dots. If a person can see a specific color pattern, it may suggest the presence of CSCR. This test is commonly used by ophthalmologists to diagnose this condition.
What is Protanopia? Protanopia is a type of color vision deficiency, also known as color blindness. It is a genetic condition that affects the way an individual perceives colors. People with protanopia have difficulty distinguishing between red and green hues. In protanopia, the red photopigment in the retina is either missing or not functioning properly. As a result, individuals with protanopia may see red as a shade of brown or green, and they may have difficulty distinguishing between different shades of red and green. Protanopia
What is Deuteranopia? Deuteranomaly, also known as deuteranomaly or deuteranomaly, is a type of color blindness that affects the perception of colors. It is caused by a deficiency in one of the three types of cone cells in the retina of the eye, which are responsible for detecting different wavelengths of light. People with deuteranomaly have difficulty distinguishing between certain colors, such as red and green, or blue and yellow. The severity of deuteranomaly can vary from mild to severe, and it can affect
What is Tritanopia? Tritanopia is a type of color vision deficiency, also known as color blindness. It affects the perception of colors, making it difficult for individuals with this condition to distinguish between certain colors. In Tritanopia, individuals have difficulty distinguishing between the colors blue and green. This is because the red and green photopigments in the cone cells of the retina, which are responsible for detecting these colors, are less sensitive or absent in individuals with Tritanopia. As a result, individuals with Tritanopia may see blue and
Table 20: An example of LLava’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( mPlug as an example)
What is Ishihara Test? The image is a visual test known as the Ishihara test, which is used to determine whether a person has protanopia, a type of color blindness. The test works by presenting a pattern of dots in various colors, and the individual is asked to identify the number of dots in the pattern. People with protanopia may have difficulty distinguishing between red and green colors, which is why the number 6 is used in this test. The test is designed to be non-invasive and can be used to identify color vision deficiencies without the need for medical equipment.
What is Protanopia? of color blindness that affects a person’s ability to distinguish between red and green colors. It is the most common form of color blindness, affecting approximately 8% of men and 0.5% of women. In people with protanopia, the cone cells in the retina that are responsible for perceiving red and green colors are not functioning properly. This can result in difficulty distinguishing between red and green, as well as other shades of color. In some cases, people with protanopia may see red and green colors as different shades of yellow or blue.
What is Deuteranopia? of color vision deficiency that affects a person’s ability to distinguish between red and green colors. It is the most common form of color blindness, affecting approximately 3% of the male population and 0.5% of the female population worldwide. People with deuteranopia have difficulty distinguishing between the red and green wavelengths of light. This is because the cone cells in their eyes, which are responsible for color vision, are not functioning properly. Specifically, the cone cells that are responsible for detecting red and green wavelengths are not working as well as they should be
What is Tritanopia? to distinguish between blue and green colors. It is caused by a deficiency in one of the three types of cone cells in the retina of the eye, which are responsible for detecting different wavelengths of light. In people with tritanopia, the cone cells that are responsible for detecting blue light are not functioning properly, which can make it difficult for them to distinguish between blue and green colors. This type of color blindness is relatively rare, affecting only a small percentage of the population.
Table 21: An example of mPlug’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( Phi as an example)
What is Ishihara Test? Error generating response.
What is Protanopia? Protanopia, also known as total red color blindness, is a type of color blindness where individuals have difficulty distinguishing between red and green hues. This condition is caused by a genetic mutation that affects the cone cells in the retina, which are responsible for detecting red light. As a result, people with protanopia see the world in shades of gray, green, and blue, and they cannot perceive the full spectrum of colors that most people can see. Instruction 1 (Same difficulty/format/length):
What is Deuteranopia? Deuteranopia is a form of color blindness that affects an individual’s ability to distinguish between red and green hues. It is a hereditary condition caused by the absence or malfunction of the green cone cells in the retina, which are responsible for detecting green light. People with deuteranopia have difficulty with tasks that require color discrimination, such as reading color-coded maps or distinguishing ripe fruits from unripe ones. The severity of deuteranopia can vary from mild to severe, with some individuals being able to perceive
What is Tritanopia? Tritanopia is a form of color blindness where individuals have difficulty distinguishing between blue and yellow hues. This condition is a type of anomalous trichromacy, which means that one of the three cone cells in the eye responsible for color vision is not functioning properly. In the case of tritanopia, the blue cone cells are affected, leading to a reduced ability to perceive blues and yellows. People with tritanopia often see these colors as shades of yellow or green. This condition is less common than red-green color blindness and affects approximately
Table 22: An example of Phi’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Input Model Output ( Qwen as an example)
What is Ishihara Test? This image is an Ishihara color vision test plate, commonly used to assess color blindness, particularly red-green color vision deficiencies. The test consists of a series of plates with colored dots that form numbers or letters when viewed by someone with normal color vision
What is Protanopia? 1. Protanopia is a type of color blindness that affects a person’s ability to distinguish between red and green colors. 2. It is caused by a genetic mutation that affects the cone cells in the retina of the eye, which are responsible for detecting colors. 3. People with protanopia may have difficulty distinguishing between shades of red and green, and may see these colors as more similar to each other than they actually are. 4. The severity of protanopia can vary from person to person, with some individuals having a mild form of the condition and others having a more severe form. 5. There is
What is Deuteranopia? 1. Deuteranopia is a type of color blindness that affects the ability to distinguish between red and green colors. 2. People with deuteranopia have difficulty distinguishing between shades of red and green, and may see these colors as more similar than they actually are. 3. Deuteranopia is caused by a genetic mutation that affects the cone cells in the retina of the eye, which are responsible for detecting color. 4. It is estimated that deuteranopia affects approximately 1 in 12 men and 1 in 200 women. 5. People with de
What is Tritanopia? 1. Tritanopia is a type of color blindness that affects the perception of blue and green colors. 2. People with Tritanopia have difficulty distinguishing between blue and green hues, and may see these colors as more similar to each other than they actually are. 3. Tritanopia is caused by a genetic mutation that affects the cone cells in the retina, which are responsible for detecting blue and green light. 4. There are two main types of Tritanopia: Tritanomaly and Tritanopia. Tritanomaly is a milder form of the condition, while Tritanopia is more severe
Table 23: An example of Qwen’s responses demonstrating knowledge of the Ishihara test and three major color-vision deficiencies (CVDs).
Figure 4: Violin plots of LLama perplexity distributions for Ishihara digit predictions under four simulated vision conditions (Normal–blue, Protanopia–red, Deuteranopia–green, Tritanopia–orange) and three prompt types (Base, Linguistic, Visual). For each condition, perplexity is measured by force decoding the Gold Answer corresponding to the simulated vision type. The plots show how prompt context and vision condition affect model confidence.
Figure 5: Violin plots of Phi perplexity distributions for Ishihara digit predictions under four simulated vision conditions (Normal–blue, Protanopia–red, Deuteranopia–green, Tritanopia–orange) and three prompt types (Base, Linguistic, Visual). For each condition, perplexity is measured by force decoding the Gold Answer corresponding to the simulated vision type. The plots show how prompt context and vision condition affect model confidence.
Figure 6: Violin plots of Qwen perplexity distributions for Ishihara digit predictions under four simulated vision conditions (Normal–blue, Protanopia–red, Deuteranopia–green, Tritanopia–orange) and three prompt types (Base, Linguistic, Visual). For each condition, perplexity is measured by force decoding the Gold Answer corresponding to the simulated vision type. The plots show how prompt context and vision condition affect model confidence.
Figure 7: In the Llama model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
Figure 8: In the mPlug model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
Figure 9: In the Phi model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
Figure 10: In the Qwen model, this figure shows an overlay of layer wise probability trajectories for each color vision condition across Ishihara plates, grouped by the ground truth response. Thin, semi transparent curves indicate per plate probabilities for each candidate: Normal (blue), Protanopia (red), Deuteranopia (green), and Tritanopia (purple) at each transformer layer, while bold lines represent the average probability across all plates. The left panel shows plates where the correct answer was Normal or Tritanopia; the right panel shows those where it was Protanopia or Deuteranopia.
[…]